### v3

# Easy conditioning far beyond Gaussian

Abstract
Estimating and sampling from conditional densities plays a critical role in statistics and data science, with a plethora of applications. Numerous methods are available ranging from simple fitting approaches to sophisticated machine learning algorithms. However, selecting from among these often involves a trade-off between conflicting objectives of efficiency, flexibility and interpretability. Starting from well known easy conditioning results in the Gaussian case, we show, thanks to results pertaining to stability by mixing and marginal transformations, that the latter carry over far beyond the Gaussian case. This enables us to flexibly model multivariate data by accommodating broad classes of multi-modal dependence structures and marginal distributions, while enjoying fast conditioning of fitted joint distributions. In applications, we primarily focus on conditioning via Gaussian versus Gaussian mixture copula models, comparing different fitting implementations for the latter. Numerical experiments with simulated and real data demonstrate the relevance of the approach for conditional sampling, evaluated using multivariate scoring rules.
Keywords: Stability by conditioning, Mixtures, Copula, Conditional sampling
1 Introduction
Modeling complex dependencies between variables is a fundamental task in statistics. Traditional methods often rely on Gaussian assumptions due to their mathematical tractability and ease of implementation. However, real-world data frequently exhibit features such as multi-modality, heavy tails, and skewness, which are not well-captured by Gaussian distributions. As a consequence, there is growing interest in more flexible approaches that can accommodate these characteristics while still permitting efficient inference and modeling.
In many applications, a comprehensive description of the stochastic behavior of some (multivariate) response variable given specific covariate values is required. Assuming a continuous framework, estimating a full conditional density allows computing a wide range of quantities of interest, such as confidence intervals or conditional quantiles. However, Conditional density estimation (CDE) is a challenging task since conditional distributions can be arbitrarily complex and subject to intricate changes throughout the input space.
Many methods have been proposed in the literature to perform the task of estimating conditional densities. Non-parametric methods for CDE such as conditional kernel density estimation [5] have been widely used but are susceptible to high-dimensional problems [14] and require careful bandwidth selection. Semi-parametric methods such as Least Square Density Ratio Estimation [25] have shown promising results in moderate dimensional problems. More flexible techniques, such as Distributional Random Forests [9], have demonstrated efficiency but they lack of explainability.
Parametric methods based on deep neural networks have also been used extensively for CDE. Mixture density networks [7] and kernel density networks [1], assume a mixture form for the conditional distribution. Normalising Flows [28] or Neural auto-regressive estimators [20] can also be used for this task. Recently, some modified versions of deep generative models allowing for conditioning such as conditional Variational Auto-Encoders (CVAE) [24], or conditional Generative Adversarial Networks (CGAN) [18] have shown efficiency for various tasks including image generation. These methods can learn complex distributions in high dimensions but require large amounts of training data and may be considered to lack interpretability.
In this work, we are investigating and presenting results on a class of generative models rooted in multivariate statistics and enjoying an elegant theoretical construction. The goal is to estimate the conditional distribution of a random vector $\mathbf{X_{1}}\in\mathbb{R}^{q}$ given an observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}\in\mathbb{R}^{p}$, based on a sample of moderate size and taking values in a continuous range, $(\mathbf{X_{1}},\mathbf{X_{2}})$ previously acquired. This setting is common in statistics, such as in probabilistic forecasting or imputation of systematically missing values, for example in medical or epidemiological applications. We aim to define an approach for conditional density estimation that is efficient, flexible, interpretable, and privacy-preserving, while also not requiring extensive training data.
Assuming conditions of existence to be fulfilled, a natural approach to estimate the conditional density $f(\mathbf{x_{1}}|\mathbf{x_{2}})$ is to first estimate the joint density by $\hat{f}(\mathbf{x_{1}},\mathbf{x_{2}})$ and then perform conditioning. Most methods have traditionally focused on the Gaussian case, leveraging the availability of analytical formulas for conditional distributions. Gaussian copula regression [21], while allowing for arbitrary continuous marginal distributions, still assumes a Gaussian dependence structure, which provides the advantage of closed-form expressions for the conditional distributions and enables rapid conditional sampling [19]. However, this reliance on a Gaussian copula imposes restrictive assumptions, particularly in its inability to capture complex (tail) dependencies and asymmetries commonly found in observational data.
By introducing the concept of trans-dimensional families of probability distributions stable by conditioning and demonstrating that this concept extends to mixtures and transformations, we show that a much broader class of probability distributions benefits from analytical conditioning. For applications, we will mainly focus on the Gaussian mixture copula model (GMCM) [27] due to its ease of implementation and its ability to capture multi-modal dependencies. Additionally, we demonstrate that incorporating automatic differentiation enhances the parameter fitting process of GMCM compared to conventional approaches for computing derivatives [26]. It is worth noting that this model can also be employed in a purely generative framework, such as for the privacy-preserving generation of synthetic data.
As a motivating example, we will consider the Wine data-set, which represents chemical analysis of wines grown in the same region in Italy. The analysis determined the quantities of 13 constituents found in each of the three types of wines. For illustrative purposes, we will focus on $2$ constituents, the Alcohol level and Malic acids concentration. Figure 1 illustrates the approach proposed using these variables. The joint density appears here to be rather well modeled by a GMCM, which conditional distribution can be calculated and simulated from efficiently as detailed throughout the paper.
Figure 1: Different steps for estimating a conditional density: from a sample of the joint distribution (left), we estimate the joint density by a GMCM (middle) and for a new value of Malics acids concentration, we estimate (or sample from) the conditional distribution of Alcohol level by using analytical formulas (right).
The remainder of the paper is organized as follows. In Section 2, we introduce the property of stability by marginalization and conditioning for families of multivariate distributions, provide some examples of families that satisfy these properties and show that it can be extended to mixtures and transformations. In Section 3, we present our method for estimating efficiently a conditional distribution by assuming a parametric form for the joint distribution and leveraging analytical formulas for conditioning. Section 4 focuses on the implementation of our algorithm for a particular case which is the GMCM. For the demonstration of the usefulness of our method, we perform numerical simulations and include real data examples in Section 5. Finally, in Section 6, we draw conclusions and outline areas for future research.
2 Stability by conditioning
2.1 Multivariate distributions stable by conditioning
2.1.1 Definition
Let us first consider a fixed dimension $d\geq 1$ and a product of spaces $\prod_{k=1}^{d}\mathcal{X}_{k}$. The $\mathcal{X}_{k}$ are assumed to be measure spaces, typically $\mathbb{R}$ or its subspaces, each endowed with the Lebesgue measure or its restrictions. We consider absolutely continuous probability distributions on $\prod_{k=1}^{d}\mathcal{X}_{k}$ with positive probability density functions $f^{(d)}(\cdot;\theta)$, where $\theta$ denotes a parameter. In this framework, we work indifferently with probability density functions (pdfs) or cumulative distribution functions (cdfs).
We denote as $\mathcal{F}^{(d)}=\{f^{(d)}(\cdot;\theta);\theta\in\Theta^{(d)}\}$ a parametric family of pdfs on $\prod_{k=1}^{d}\mathcal{X}_{k}$ with parameters belonging to a set $\Theta^{(d)}$.
We define a trans-dimensional family of pdfs by $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $1\leq D\leq+\infty$. The reason why we are concerned with this notion is because we will question the stability of such families by marginalization and conditioning with respect to some components.
Let us now introduce sets of indices that will be used for marginal distributions. For a given dimension $d\geq 1$, the sets of $\ell$ strictly ordered indices are:
$I_{\ell}^{(d)}=\{\mathbf{i}=(i_{1},\ldots,i_{\ell}),~{}1\leq i_{1}<\ldots<i_{\ell}\leq d\},~{}\text{with}~{}1\leq\ \ell\leq d.$
For conditional distributions, we define a set of pairs of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ with elements $\mathbf{i}$ and $\mathbf{j}$ satisfying:
  • $\mathbf{i}=(i_{1},\ldots,i_{\ell})$ set of $\ell$ indices with $1\leq i_{1}<\ldots<i_{\ell}\leq d$, $11\leq\ell\leq d-$ 1,
  • $\mathbf{j}=(j_{1},\ldots,j_{m})$ set of $m$ indices with $1\leq j_{1}<\ldots<j_{m}\leq d$, $1\leq m\leq d-\ell$,
  • $\text{set}(\mathbf{i})\cap\text{set}(\mathbf{j})=\emptyset$ where the set of a vector is the set of all its components.
For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$ and a set of $\ell$ strictly ordered indices $\mathbf{i}\in I_{\ell}^{(d)}$, we define the function:
[…]
which applies the marginalization of a $d$-dimensional density over $\ell$ variables.
To extend this operation to a trans-dimensional setting, we extend the marginalization operator as follows. For $d\geq 1$ and $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we define :
$\text{Marg}(f^{(d)}(\cdot;\theta))=\bigcup\limits_{\ell=1}^{d-1}\{f_{\mathbf{i}}(\cdot;\theta),~{}\mathbf{i}\in I_{\ell}^{(d)}\}$
Now if we apply Marg to the entire family of distributions $\mathcal{F}^{(d)}$, we get:
$\text{Marg}(\mathcal{F}^{(d)})=\left\{\text{Marg}(f^{(d)}(\cdot;\theta)),\theta\in\Theta^{(d)}\right\}$
and finally for a trans-dimensional family of distributions $\mathcal{F}$ we set:
$\text{Marg}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\text{Marg}(\mathcal{F}^{(d)})$
A trans-dimensional family of probability distributions $\mathcal{F}$ is said to be stable by marginalization if $\text{Marg}(\mathcal{F})\subset\mathcal{F}$.
Let us know consider conditioning. Since the pdfs considered here are positive, all conditional distributions are defined. For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$, a pair of sets of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and observed components $(x_{j_{1}},\ldots,x_{j_{m}})$, we define:
$f_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}})):(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{i_{k}}\rightarrow\frac{f_{\left[\mathbf{i},\mathbf{j}\right]}(x_{i_{1}},\ldots,x_{i_{\ell}},x_{j_{1}},\ldots,x_{j_{m}})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}$
where $\left[\mathbf{i},\mathbf{j}\right]=(i_{1},\ldots,i_{\ell},j_{1},\ldots,j_{m})$.
In the same manner as above for marginalization, we define a conditioning operator taking a pdf in a parametric family as input and returning every possible conditional densities i.e for all pairs of sets of strictly ordered indices and every possible set of conditioning variables. For a function $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we have :
$\text{Cond}(f^{(d)}(\cdot,\theta))=\left\{f^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}});\theta),~{}(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d},~{}(x_{j_{1}},\ldots,x_{j_{m}})\in\prod_{k=1}^{m}\mathcal{X}_{i_{k}}\right\}$
By applying it to the trans-dimensional family of pdf $\mathcal{F}$ we get:
$\text{Cond}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\left\{\text{Cond}(f^{(d)}(\cdot;\theta)),~{}\theta\in\Theta^{(d)}\right\}$
A trans-dimensional family of probability distributions $\mathcal{F}$ is said to be stable by conditioning if $\text{Cond}(\mathcal{F})\subset\mathcal{F}$.
2.1.2 Examples
Stability by marginalization and conditioning can be demonstrated by showing that the marginal or conditional distributions remain within the same family of distributions, albeit in a lower dimension. Here, we will present several examples of distribution families that exhibit this property.
  • Gaussian distribution: For a fixed dimension $d\geq 1$, we consider
    $\mathcal{F}^{(d)}=\left\{f^{(d)}(x_{1},\ldots,x_{d};\mu,\Sigma)=\frac{\exp(-\frac{1}{2}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu))}{\sqrt{(2\pi)^{d}|\Sigma|}},\theta=(\mu,\Sigma)\in\Theta^{(d)}\right\}$
    where $\Theta^{(d)}=\mathbb{R}^{d}\times\text{PD}_{d}$ where $\text{PD}_{d}$ is the set of all symmetric positive definite matrix on $\mathbb{R}^{d\times d}$. The trans-dimensional family of Gaussian distributions is defined by: $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D=+\infty$. By properties of the multivariate Gaussian distribution, we know that the marginal and conditional distributions remains Gaussian with parameters that can be computed analytically via well-known formulas. We have that $\text{Marg}(\mathcal{F})\subset\mathcal{F}$ and $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, the trans-dimensional family of Gaussian distributions is stable by marginalization and by conditioning.
  • Student $t$-distribution: Another example is the family of multivariate Student t-distributions, which is a generalization for multivariate normal distributions allowing for heavier tails. Unlike the multivariate Gaussian distribution, it can model tail dependencies and it includes an additional “degrees of freedom” parameter, which controls the shape of the distribution, with higher values leading to distributions that resemble the multivariate Gaussian, while lower values result in heavier tails.
    It is defined as a scaled mixture of Gaussian distributions. In particular, if $\textbf{Y}\sim\mathcal{N}_{d}(0,\Sigma)$ and $u\sim\chi^{2}_{\nu}$ are independent then $\textbf{X}=\frac{\textbf{Y}}{\sqrt{\frac{u}{\nu}}}+\mu$ is such that $\textbf{X}\sim t_{d}(\mu,\Sigma,\nu)$, where $\chi^{2}_{\nu}$ represents a chi-square distribution with $\nu$ degrees of freedom. We say that X follows a multivariate $t$-distribution with mean $\mu$, covariance matrix $\Sigma$ and $\nu$ degrees of freedom.
    For $d\geq 1$, we denote the family of multivariate Student $t$-distribution as:
    $\mathcal{F}^{(d)}=\left\{f^{(d)}(\mathbf{x};\mu,\Sigma,\nu),\theta=(\mu,\Sigma,\nu)\in\Theta^{(d)}=\mathbb{R}^{d}\times\text{PD}_{d}\times\left[0,+\infty\right[\right\}$
    with $f^{(d)}(\mathbf{x};\mu,\Sigma,\nu)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})(\nu\pi)^{\frac{d}{2}}|\Sigma|^{\frac{1}{2}}}(1+\nu^{-1}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu))^{-\frac{(\nu+d)}{2}}$.
    The multivariate Student $t$-distribution is stable by marginalization and conditioning [11]; more details are included in the appendix.
  • Other examples: More generally the family of elliptically contoured distributions also enjoys these properties [8]. Other examples of families of multivariate distributions stable by marginalization and conditioning are the multivariate unified skew normal and skew Student $t$-distributions [3]. Unlike elliptical distributions, they allow for skewness.
  • Counter example: Let us now consider a simple counter-example of a family of multivariate distributions which is stable by marginalization but not stable by conditioning. For a given $d\geq 1$, consider the family of multivariate Student $t$-distributions with a fixed degrees of freedom $\nu\in\left[0,+\infty\right[$ i.e
    $\mathcal{F}^{(d)}_{\nu}=\left\{f^{(d)}(\mathbf{x};\mu,\Sigma),\theta=(\mu,\Sigma)\in\Theta^{(d)}_{\nu}=\mathbb{R}^{d}\times\text{PD}_{d}\right\}$
    where $f^{(d)}(\mathbf{x};\mu,\Sigma)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})(\nu\pi)^{\frac{d}{2}}|\Sigma|^{\frac{1}{2}}}(1+\nu^{-1}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu))^{-\frac{(\nu+d)}{2}}$ and the trans-dimensional family is defined by $\mathcal{F}_{\nu}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}_{\nu}$ with $D=+\infty$. Then by properties of the multivariate Student $t$-distribution [11], this family is stable by marginalization but not by conditioning, because when conditioning, we change the degrees of freedom of the distribution.
2.2 Main properties
2.2.1 Extension to finite mixtures
Let us first define a “mixing” operator which takes as input a parameterized family of multivariate pdfs with a fixed dimension $d\geq 1$ and returns the set of probabilities densities that are finite mixtures of elements of this family.
Consider again a parametrized family $\mathcal{F}^{(d)}=\{f^{(d)}(\cdot;\theta);\Theta\in\Theta^{(d)}\}$ of pdfs. We define the mixing operator as:
$\text{Mix}(\mathcal{F}^{(d)})=\left\{\sum_{k=1}^{K}\alpha_{k}f^{(d)}(\cdot;\theta_{k}),~{}K\geq 1,~{}\theta_{1},\ldots,\theta_{K}\in\Theta^{(d)},~{}\alpha_{1},\ldots,\alpha_{K}\geq 0~{}\text{s.t}~{}\sum_{k=1}^{K}\alpha_{k}=1\right\}$
For a trans-dimensional family of distributions $\mathcal{F}$, we define its mixing set of distributions as the union over the mixing sets of the families with fixed dimensions i.e
$\text{Mix}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\text{Mix}(\mathcal{F}^{(d)})$
We now present the crucial results that extend the concepts of stability by marginalization and conditioning to finite mixtures.
Theorem 2.1.
If a parametrized family of trans-dimensional probability distributions $\mathcal{F}$ is stable by conditioning (resp. marginalization) then $\text{Mix}(\mathcal{F})$ is also stable by conditioning (resp. marginalization).
Proof: Let consider $f$ in $\text{Mix}(\mathcal{F})$. There exists $d\in\{1,\ldots,D\}$, $K\geq 1$, $\theta_{1},\ldots,\theta_{K}\in\Theta^{(d)}$, $\alpha_{1},\ldots,\alpha_{K}\geq 0~{}\text{s.t}~{}\sum_{k=1}^{K}\alpha_{k}=1$ such that:
$f(\cdot)=\sum_{k=1}^{K}\alpha_{k}f^{(d)}(\cdot;\theta_{k})$
For every pair of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and conditioning variables $(x_{j_{1}},\ldots,x_{j_{m}})\in\prod_{k=1}^{m}\mathcal{X}_{j_{k}}$, we have that $\forall(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{j_{k}}$:
$\begin{split}f_{\mathbf{i}|\mathbf{j}}((x_{i_{1}},\ldots,x_{i_{\ell}})|(x_{j_{1}},\ldots,x_{j_{m}}))&=\frac{f_{[\mathbf{i},\mathbf{j}]}(x_{i_{1}},\ldots,x_{i_{\ell}},x_{j_{1}},\ldots,x_{j_{m}})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}\\ &=\frac{1}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}\sum_{k=1}^{K}\alpha_{k}f^{(d)}_{[\mathbf{i},\mathbf{j}]}(x_{i_{1}},\ldots,x_{i_{\ell}},x_{j_{1}},\ldots,x_{j_{m}};\theta_{k})\\ &=\sum_{k=1}^{K}\frac{\alpha_{k}f^{(d)}_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}};\theta_{k})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}f^{(d)}_{\mathbf{i}|\mathbf{j}}((x_{i_{1}},\ldots,x_{i_{\ell}})|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\\ &=\sum_{k=1}^{K}\tilde{\alpha}_{k}(x_{j_{1}},\ldots,x_{j_{m}})f^{(d)}_{\mathbf{i}|\mathbf{j}}((x_{i_{1}},\ldots,x_{i_{\ell}})|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\end{split}$
where $\tilde{\alpha_{k}}(x_{j_{1}},\ldots,x_{j_{m}})=\frac{\alpha_{k}f^{(d)}_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}};\theta_{k})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}$ which depends only on indices in $\mathbf{j}$.
Since $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, then for every $k\in\{1,\ldots,K\},~{}f^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\in\mathcal{F}$ and the conditional distribution $f_{\mathbf{i}|\mathbf{j}}$ is a mixture of elements of $\mathcal{F}$ with weights $\tilde{\alpha}_{k}$ so we can conclude that $f_{\mathbf{i}|\mathbf{j}}\in\text{Mix}(\mathcal{F})$. The proof for marginalization is analogous. $\square$
This result can similarly be applied to cdfs. Note that if we can compute analytically the conditioned mixing weights $\tilde{\alpha}_{k}$ and if we can sample from the conditional distributions of each component $f_{\mathbf{i}|\mathbf{j}}(\cdot;\theta_{k})$, we can produce conditional samples of a mixture of distributions.
2.2.2 Extension to transformations
We define, similarly as in the previous section for pdfs, parametric families of cdfs on $\prod_{k=1}^{d}\mathcal{X}_{k}$ as $\mathcal{F}^{d}=\left\{F^{(d)}(\cdot;\theta);\theta\in\Theta^{(d)}\right\}=\left\{F^{(d)}_{\theta};\theta\in\Theta^{(d)}\right\}$ and trans-dimensional families as $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D\leq+\infty$.
For fixed $d\geq 1$ and $\mathcal{X}^{\prime}_{i},~{}i=1,\ldots,d$ measure spaces, we define the class of functions:
$\Xi^{d}=\left\{\xi=\begin{pmatrix}\xi_{1}\\ \vdots\\ \xi_{d}\end{pmatrix}~{}\text{s.t}~{}\xi_{i}\in\mathcal{C}^{1}_{\uparrow}(\mathcal{X}_{i},\mathcal{X}^{\prime}_{i}),~{}i=1,\ldots,d\right\}$
where $\mathcal{C}^{1}_{\uparrow}(\mathcal{X}_{i},\mathcal{X^{\prime}}_{i})$ is the class of increasing $\mathcal{C}^{1}$ diffeomorphisms from $\mathcal{X}_{i}$ to $\mathcal{X}^{\prime}_{i}$. We introduce a “push-forward” operator, which takes as input an element of a parametric family of cdfs and returns every possible cdfs obtained by applying marginal $\mathcal{C}^{1}_{\uparrow}$ diffeomorphisms to the components, by:
$\Xi^{d}\#F^{(d)}_{\theta}=\left\{F^{(d)}_{\theta}\circ\xi^{-1};\xi\in\Xi^{d}\right\}$
Note that elements of $\Xi^{(d)}\#F^{(d)}_{\theta}$ are probability distributions on $\prod_{i=1}^{d}\mathcal{X^{\prime}}_{i}$ and that we restricted to differentiable bijections such that the obtained probability distributions are still absolutely continuous.
By applying the operator to the whole family of cdfs $\mathcal{F}^{(d)}$ we get:
$\Xi^{d}\#\mathcal{F}^{(d)}=\left\{\Xi^{d}\#F^{(d)}_{\theta};\theta\in\Theta^{(d)}\right\}$
If we define the class of functions with increasing marginal $\mathcal{C}^{1}$ diffeomorphisms as
$\Xi=\bigcup\limits_{d=1}^{D}\Xi^{d}$
we can now apply the "push-forward" operator to a trans-dimensional family of cdfs $\mathcal{F}$ by:
$\Xi\#\mathcal{F}=\bigcup\limits_{d=1}^{D}\Xi^{d}\#\mathcal{F}^{(d)}$
We call the family of probability distributions $\Xi\#\mathcal{F}$ the push-forward family of $\mathcal{F}$-distributions. In the next section, we will relate this concept to copulas and meta families of distributions. At this point, we present the results that extend the stability of marginalization and conditioning to transformations.
Theorem 2.2.
If a parametrized family of trans-dimensional probability distributions $\mathcal{F}$ is stable by conditioning (resp. marginalization) then $\Xi\#\mathcal{F}$ is also stable by conditioning (resp. marginalization).
Proof: Let $F\in\Xi\#\mathcal{F}$. There exists $d\in\{1,\ldots,D\},~{}\theta\in\Theta^{(d)},~{}\xi\in\Xi^{d}$ such that:
$\forall(x_{1},\ldots,x_{d})\in\prod_{k=1}^{d}\mathcal{X}_{k},~{}F(x_{1},\ldots,x_{d})=(F^{(d)}\circ\xi^{-1})(x_{1},\ldots,x_{d};\theta)=F^{(d)}(\xi_{1}^{-1}(x_{1}),\ldots,\xi_{d}^{-1}(x_{d});\theta)$
For every pair of strictly of ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and conditioning variables $(x_{j_{1}},\ldots,x_{j_{m}})\in\prod_{k=1}^{m}\mathcal{X}_{j_{k}}$, we have that $\forall(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{j_{k}}$:
$\begin{split}F_{\mathbf{i}|\mathbf{j}}(x_{i_{1}},\ldots,x_{i_{\ell}}|x_{j_{1}},\ldots,x_{j_{m}})&=F^{(d)}_{\mathbf{i}|\mathbf{j}}(\xi_{i_{1}}^{-1}(x_{i_{1}}),\ldots,\xi_{i_{\ell}}^{-1}(x_{i_{\ell}})|\xi_{j_{1}}^{-1}(x_{j_{1}}),\ldots,\xi_{j_{m}}^{-1}(x_{j_{m}});\theta)\\ &=F^{(d)}_{\mathbf{i}|\mathbf{j}}(z_{i_{1}},\ldots,z_{i_{\ell}}|z_{j_{1}},\ldots,z_{j_{m}};\theta)\end{split}$
with $z_{i_{k}}=\xi^{-1}_{i_{k}}(x_{i_{k}}),k=1,\ldots,\ell$ and $z_{j_{k}}=\xi^{-1}_{j_{k}}(x_{j_{k}}),k=1,\ldots,m$. Since $\mathcal{F}$ is stable by conditioning, then $F^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot;\theta)$ belongs to to $\mathcal{F}$. The proof is similar for marginalization. $\square$
Note that we have only considered univariate transformations. However, for a fixed set of conditioning indices, it is worth mentioning that by following the same reasoning as in the proof, we can extend the property to blockwise bijective transformations. This goes beyond the scope of the present paper.
3 Methodology for Estimating Conditional Distributions
As highlighted in the introduction, our focus is on estimating the conditional distributions of a (multivariate) response variable. This section delineates our methodology.
Assume we want to condition on an given set of variables. Without loss of generality, we can rearrange the components of a random variable $\mathbf{X}$ in two groups of variables $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2})$ such that the first group $\mathbf{X_{1}}$ contains the variables that we want to estimate conditional on the second group of variables $\mathbf{X}_{2}$. By assuming that we can fit our models on a previously acquired sample $(\mathbf{X_{1}},\mathbf{X_{2}})$ taking values on a continuous scale, and given a new an observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}$, we derive a procedure to estimate the conditional distribution of $\mathbf{X}_{1}$ given that $\mathbf{X}_{2}=\mathbf{x}_{2}$.
We consider the scenario of real-valued random variables, corresponding to $\prod_{k=1}^{d}\mathcal{X}_{k}=\mathbb{R}^{d}$ and we focus exclusively on absolutely continuous random variables.
Our generative procedure to estimate a conditional distribution can be divided in two main steps:
  • estimate the joint distribution by a parametric form.
  • perform conditioning of the estimated joint by using analytical formulas.
In modeling the joint distribution, much of the existing literature focuses on the Gaussian case due to the simplicity of its analytical formula for conditional distributions. However, multivariate Gaussian distributions have notable limitations, as they often fail to adequately model skewed, heavy-tailed, or multi-modal data.
Alternative multivariate distributions, such as Student’s t-distributions [11] or skew-normal distributions [3], offer more expressiveness and are stable by conditioning, see appendix for details. Yet, these distributions also fall short in scenarios requiring the modeling of multi-modal data and are confined to have marginal distributions of the same family along each axis.
By leveraging the extension of stability by conditioning to mixtures and transformations discussed in the preceding section, we achieve considerably greater flexibility in modeling the joint distribution. The extension to transformations suggests that we can identify a latent space where stability by conditioning holds. To develop a fast and generic approach for estimating these conditional distributions, our interest lies particularly in families of distributions for which conditional distributions are known in closed form.
By using univariate transformations, we can model distributions with arbitrary marginals. Specifically, if we consider the univariate transformations to be the marginal cumulative distribution functions, we obtain a space with uniform marginals and can subsequently apply copula theory, as detailed in the next section.
3.1 Estimate the joint multivariate distribution with copulas
From a $d$-dimensional sample $\mathcal{D}=(x_{1i},\ldots,x_{di})_{i=1}^{n}$ of independently drawn realizations, we want to estimate the distribution of the underlying random vector $\mathbf{X}$ taking values in $\mathbb{R}^{d}$. We will use a two-steps procedure based on copulas.
Copulas provides a versatile framework for multivariate analysis. Following Sklar’s theorem [22], for any distribution function $F$ with continuous margins $F_{j},~{}j=1,\ldots,d$, there exists a unique copula $C$ such that
$F(x_{1},\ldots,x_{d})=C(F_{1}(x_{1}),\ldots,F_{d}(x_{d})),~{}\forall\textbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$
A multivariate distribution function can then be modeled separately by its marginals and its copula where the copula characterizes the dependence structure between the variables.
There exist several parametric families of copulas. In particular, let’s introduce the notion of “implicit copula” which refers to the copula implicit in the multivariate distribution of a continuous random vector.
By “inverting” Sklar’s theorem, if we assume that the marginal distributions $F_{j},~{}j=1,\ldots,d$ are strictly increasing and continuous, we obtain that the copula can be written:
$C(u_{1},\ldots,u_{d})=F(F_{1}^{-1}(u_{1}),\ldots,F^{-1}_{d}(u_{d})),~{}0\leq u_{1},\ldots,u_{d}\leq 1$
For example, the Gaussian copula, which is the copula implicit in a multivariate Gaussian distribution can be expressed:
$C_{P}(\mathbf{u})=\Phi_{P}(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d}))$
where $\Phi$ is the cdf of the univariate standard normal distribution and $\Phi_{P}$ is the cdf of the multivariate $\mathcal{N}_{d}(\mathbf{0},\mathbf{P})$ distribution with $\mathbf{P}$ a correlation matrix.
Differentiating the general formula with respect to $\mathbf{u}$, we get the density of the implicit copula by:
$c(u_{1},\ldots,u_{d})=\frac{\partial^{d}}{\partial u_{1}\ldots\partial u_{d}}C(u_{1},\ldots,u_{d})=\frac{f(z_{1},\ldots,z_{d})}{\prod_{i=1}^{d}f_{i}(z_{i})}$
where $z_{j}=F_{j}^{-1}(u_{j}),~{}j\in\{1,\ldots,d\}$, we say that $\mathbf{z}=(z_{1},\ldots,z_{d})$ is the quantile transform of $\mathbf{u}$.
If the marginals are known, we can transform the observations $\mathbf{x}_{i}$ to pseudo-observations $\mathbf{u}_{i}$, with uniform marginals, by using the Probability Integral Transform (PIT), $u_{ji}=F_{j}(x_{ji}),~{}\forall j\in\{1,\ldots,d\}$. And then to fit a parametric copula $c(u_{1},\ldots,u_{d};\theta)$ to the pseudo-observations by estimating the copula parameters $\theta$ with maximum-likelihood.
The optimization problem is then:
$\hat{\theta}\in\underset{\theta\in\Theta}{\mathrm{argmin}}\prod_{i=1}^{n}c(u_{1i},\ldots,u_{di};\theta)$
However, in practice we don’t know the marginal distributions and we need to estimate them from the data. Parametric or non-parametric methods can be used to estimate the marginal distributions. Parametric estimation of the marginals is used in the inference for margins (IFM) method [15], while non-parametric estimation using the empirical distribution function is part of the canonical maximum likelihood estimation (CMLE) method [13], which has been proven to be asymptotically consistent.
Once the marginals have been estimated, then we can plug-in the “estimated pseudo observations” , $\hat{u}_{ji}=\hat{F}_{j}(x_{ji})$ in the maximum-likelihood optimization problem.
A $d$-variate meta distribution is constructed by linking a $d$-variate implicit copula with $d$ arbitrary marginal cdfs. For example, a meta Gaussian distribution has a Gaussian copula with arbitrary marginals. Meta elliptical distributions have been studied [12].
More formally, for a family of distributions $\mathcal{F}^{(d)}$ in dimension $d$ such as Gaussian or Student $t$-distributions we define its meta family as :
$\text{Meta}(\mathcal{F}^{(d)})=\left\{F(\cdot)=(C_{\theta}\circ\tilde{F})(\cdot),~{}\theta\in\Theta^{(d)},~{}\tilde{F}\in\mathcal{M}^{d}\right\}$
where $\left\{C_{\theta},~{}\theta\in\Theta^{(d)}\right\}$ is a family of parametric copulas in dimension $d$ such that $\forall\theta\in\Theta^{(d)},~{}F_{\theta}\in\mathcal{F}^{(d)}~{}\text{has copula}~{}C_{\theta}$ and :
$\mathcal{M}^{d}=\left\{F=\begin{pmatrix}F_{1}\\ \vdots\\ F_{d}\end{pmatrix},~{}F_{1},\ldots,F_{d}~{}\text{absolutely continuous and strictly increasing cdfs}\right\}$
And for a trans-dimensional family $\mathcal{F}$ defined as in the last section, we get:
$\text{Meta}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\text{Meta}(\mathcal{F}^{(d)})$
We can remark that for a given family $\mathcal{F}$ its push-forward family includes its meta family. The proof showing that stability by conditioning of a given trans-dimensional family $\mathcal{F}$ implies the stability by conditioning of its meta family, closely resembles that of Theorem 2.2.
In our procedure, we assume that the joint distribution of the data is a meta distribution with a parametric copula from a well known family. In particular, we choose families of distributions for which are stable by conditioning. For example, the conditional distributions of a meta Gaussian distributions are also meta Gaussian.
We should mention that the stability of meta distributions corresponds to the stability of copulas. Our research focuses on parametric families of copulas for which the conditional copulas belongs to the same parametric family.
3.2 Conditioning in a latent space
Once we have estimated the joint distribution, the second step of our method is to estimate the conditional distribution (get conditional samples) given a new observation of the covariates. Since we assume that the data follow a "meta stable" distribution in order to perform exact conditioning, we will be able to find a latent space in which we have stability by conditioning. In practice, we will focus on distributions for which there is an analytical formula for the conditional distributions.
Copulas transform the marginals to the uniform space $\mathcal{U}$ because it is usually easier to capture multivariate dependence using $C$ on $[0,1]^{d}$. We go one step further by transforming the marginals to a latent space $\mathcal{Z}$ [23], in which the marginals are the margins of an implicit multivariate distribution driving the copula. The two consecutive transformations are explained in the following scheme:
where $\Psi_{i}(\cdot;\hat{\theta})$ represents the cumulative distribution function of the $i$-th marginal of an implicit multivariate distribution linked to the copula parameters $\hat{\theta}$ estimated in the last section. In the latent space $\mathcal{Z}$, the samples follow this implicit multivariate distribution.
By our assumption of "meta stable" distribution, we can use analytical formulas for conditioning in the latent space. Finally we transform back to the original space to get an estimate of the conditional distribution with the right marginals.
The full pseudo-algorithm to estimate a conditional distribution function is presented next:
Algorithm 1 Estimate the cdf of a conditional distribution
0:  $\mathcal{D}=(x_{1i},\ldots,x_{di})_{i=1}^{n}$ a sample from $\mathbf{X}\in\mathbb{R}^{d}$, $\ell$ the dimension of the response, a new point $\mathbf{x}^{(new)}_{\ell+1:d}=(x^{(new)}_{\ell+1},\ldots,x^{(new)}_{d})$.
  • Estimate the marginal distributions $F_{j},~{}j\in\{1,\ldots,d\}$ by $\hat{F}_{j}$.
  • Transform each marginal to approximately uniform by PIT, $\hat{u}_{j}=\hat{F}_{j}(x_{j})$.
  • Learn the copula function, which represents dependence structure by assuming a parametric form $c(\hat{u}_{1},\ldots,\hat{u}_{d},\theta)$. This is the implicit copula of a multivariate distribution with marginals $\Psi_{1},\ldots,\Psi_{d}$.
  • Transform the new observed point to the latent space $\mathcal{Z}$ by applying the transform
    $\mathbf{z}^{(new)}_{\ell+1:d}=(z^{(new)}_{\ell+1},\ldots,z^{(new)}_{d})=(\Psi^{-1}_{\ell+1}\circ\hat{F}_{\ell+1}(x_{\ell+1}^{(new)}),\ldots,\Psi^{-1}_{d}\circ\hat{F}_{d}(x_{d}^{(new)}))$
  • Perform conditioning in the latent space by using the analytical formulas to get an estimate of the cdf:
    $\hat{F}_{\mathbf{Z}_{1}|\mathbf{Z}_{2}=z^{(new)}_{\ell+1:d}}(z_{1}^{(new)},\ldots,z_{\ell}^{(new)}|z^{(new)}_{\ell+1:d})$
  • Transform back to get an estimate of the cdf in the original space by:
    $\hat{F}_{\mathbf{X}_{1}|\mathbf{X}_{2}=\mathbf{x}^{(new)}}(x_{1},\ldots,x_{\ell}|\mathbf{z}_{\ell+1:d}^{(new)})=\hat{F}_{\mathbf{Z}_{1}|\mathbf{Z}_{2}=\mathbf{z}^{(new)}_{\ell+1:d}}(\hat{F}_{1}^{-1}\circ\Psi_{1}(z_{1}^{(new)}),\ldots,\hat{F}_{\ell}^{-1}\circ\Psi_{\ell}(z_{\ell}^{(new)})|\mathbf{z}_{\ell+1:d}^{(new)})$
The algorithm above can be easily adapted to sample from the conditional distribution. The full workflow for estimating conditional distributions is illustrated on a $2$-dimensional example in figure 2.
Figure 2: Workflow of the conditioning algorithm on a $2$-dimensional example.
4 Implementation of the method
Our method’s versatility stems from its ability to choose any dependence structure among families of copulas stable by conditioning, which includes mixtures, and to model the joint distribution with arbitrary marginals. The primary challenge in our procedure lies in fitting a copula in the uniform space. This involves balancing the complexity of the dependence structure against the precision with which we can estimate the copula parameters. For this paper, we have opted to use a Gaussian mixture copula with a limited number of components. However, alternative dependence structures such as Student’s $t$-mixture copulas or Skew-Normal mixture copulas represent exciting avenues for future research.
4.1 Gaussian Mixture Copula Model
The Gaussian mixture copula [27] is the copula implicit in a Gaussian mixture model. It allows to capture multi-modality in the dependence structure. The density of a Gaussian mixture model (GMM) with $K$ components is given by:
$\psi(x_{1},\ldots,x_{d},\theta)=\sum_{k=1}^{K}\alpha_{j}\phi(x_{1},\ldots,x_{d};\mu_{k},\Sigma_{k})$
with $\alpha_{k}\geq 0,~{}\forall k\in\{1,\ldots,K\}$ such that $\sum_{k=1}^{K}\alpha_{k}=1$, the mixing weights and $\mu^{(k)},~{}\Sigma^{(k)}$ the mean and covariance matrix of the $k$-th component. $\theta=\{\alpha_{i},\mu_{i},\Sigma_{i}\}_{i=1}^{K}$ is the set of all parameters and $\phi$ the pdf of a Gaussian distribution.
Then from the equation 3.1, we have that the density of a Gaussian mixture copula with parameters $\theta$ is given by:
$c_{gmc}(u_{1},\ldots,u_{d};\theta)=\frac{\psi(\Psi^{-1}_{1}(u_{1}),\ldots,\Psi^{-1}_{d}(u_{d}))}{\prod_{i=1}^{d}\psi_{i}(\Psi_{i}^{-1}(u_{i}))},~{}0\leq u_{1},\ldots,u_{d}\leq 1$
where $\psi_{j}$ and $\Psi_{j}^{-1}$ denote the marginal density and inverse distribution function of the GMM along the $j$-th dimension .
Similarly, we can define the Student $t$-mixture copula [17] or the unified Skew-Normal mixture copula as others implicit copulas.
4.1.1 Fitting of a GMCM
Estimating parameters of a GMCM is a challenging task for several reasons as it has been discussed in [26].
The first challenge is that copulas are invariant under marginal increasing transformations. Specifically, if $\mathbf{X}=(X_{1},\ldots,X_{d})^{t}$ is a random vector with continuous marginal distributions and copula $C$, and $T_{1},\ldots,T_{d}$ are strictly increasing functions. Then $(T_{1}(X_{1}),\ldots,T_{d}(X_{d}))^{t}$ also has copula $C$.
In the case of a GMCM, this property implies a lack of identifiability of the parameters, which makes the fitting of the parameters more difficult. This can be seen through the following proposition [26]:
Proposition 4.1.
(Non identifiability of GMCM)
Let $\theta=\{\alpha_{i},\mu_{i},\Sigma_{i}\}_{i=1}^{K}$ and $\theta^{\prime}=\{\alpha_{i},A\mu_{i}+\mathbf{b},A^{t}\Sigma_{i}A\}_{i=1}^{K}$ two valid set of parameters with $\mathbf{b}\in\mathbb{R}^{d},~{}A\in\mathbb{R}^{d\times d}$. Then the likelihood of the two Gaussian mixture copula models with these parameters are the same i.e $\forall\mathbf{u}\in\mathbb{R}^{d},~{}c_{gmc}(u_{1},\ldots,u_{d};\theta)=c_{gmc}(u_{1},\ldots,u_{d};\theta^{\prime})$
This lack of identifiability results in the likelihood of a GMCM having infinitely many maximizers. To address this issue, Bilgrau [6] proposed standardizing the first component to have zero mean and unit variance along each dimension, which resolves the identifiability issue in most cases.
Another challenge in fitting GMCM is the high computational cost associated with functional evaluations, as the quantile function of a Gaussian mixture model isn’t known in closed form. To address this, we can use numerical approximation schemes leveraging the monotonicity of the cdf, although they tend to be computationally expensive. Alternatively, methods based on the bisection algorithm, such as the one developed by Chandrupatla [10], can be employed to find the inverse of a cdf. Since the derivatives of the Gaussian mixture model’s cdf are known analytically, a Newton-type algorithm can also be effectively used in this context.
In the literature, two primary methods have been proposed for estimating the parameters of a GMCM via direct gradient-based optimization. The classical approach involves using finite difference methods (denoted as FD in the rest of this paper) to estimate the derivatives. This approach approximates the derivatives numerically, which can be computationally expensive and less precise.
Recently, a more efficient method, leveraging modern probabilistic programming language tools such as automatic differentiation (denoted as AD), has been developed. Automatic differentiation computes the exact values of the derivatives by leveraging the chain rule, significantly enhancing both computational efficiency and accuracy.
Another possible approach for parameter estimation is using an Expectation-Maximization (EM) type algorithm. However, due to the changing inverse distribution functions along the margins with every parameter update, the inputs of the EM algorithm are not fixed, rendering the traditional EM algorithm unsuitable. Although a pseudo-EM algorithm (denoted as PEM) has been proposed and discussed in the literature, it lacks convergence guarantees and often shows suboptimal performance.
The methods FD and PEM are implemented in the R package GMCM [6]. Meanwhile, the AD method has been developed in Python by Tewari.
4.1.2 Comparison of the copula fitting methods
To compare the different optimization methods described above (AD, FD and PEM), we simulated data following a GMCM with known parameters and attempted to recover its values using different fitting procedures. Because GMCM parameters are not identifiable, we cannot directly compare the recovered parameter values. Instead, we need to compare metrics such as the log-likelihood or distances to the true distribution. We used the Cramér distance to quantify the differences between empirical multivariate distributions. We compared samples of the true underlying distributions to samples of the distribution with the fitted parameters and repeated the experiments several times to take into account the uncertainty in the sampling of a GMCM. In the results, we indicated the average Cramér distance and average log-likelihood as well as the standard deviations.
To ensure generality, we simulated data from several parameter configurations with varying dimensions and numbers of components, see details in the appendix.
Table 1: Fitting performances of the different optimization methods for a $2D$ example in terms of distance to a sample from the true underlying distribution.
Log-likelihood Cramer
Mean SD Mean SD
AD 208.7 16.9 0.278 $8.4\times 10^{-5}$
PEM -151.7 34.2 0.422 $9.1\times 10^{-5}$
FD -154.4 35.3 0.406 $8.8\times 10^{-5}$
Table 2: Fitting performances of the different optimization methods for a $3D$ example in terms of distance to a sample from the true underlying distribution.
Log-likelihood Cramer
Mean SD Mean SD
AD 1104.2 44.1 0.330 $8.9\times 10^{-5}$
PEM 1069.2 35.7 0.360 $8.80\times 10^{-5}$
FD 1063.8 43.7 0.364 $9.3\times 10^{-5}$
From the results above, it seems that the direct optimization of the likelihood function by using automatic differentiation tools (AD) performs better than the other methods for several parameter values. That is why we rely on this method for GMCM fitting in the experiments presented next.
5 Numerical experiments
We split our data set randomly into a training set containing $80\%$ of the observations and a test set containing the remaining $20\%$ of the observations. We use the training set to learn the joint multivariate distribution and we sample $1000$ points from the estimated conditional distribution at every test points.
We then evaluate the performance of our approach comparing the conditional samples to the observed values using scoring rules. In the univariate case, we use the Continuous Ranked Probability Score (CRPS) and the logarithmic score. For multivariate outcomes, we employ the Energy score and the Variogram score [16], with order $r=0.5$. In particular, we will examine the mean value of the various scores across all test points, averaged over multiple train/test splits.
We denote the method presented in section 4 as GMCM. The marginal distributions are learned parametrically by assuming a Gaussian mixture model, with the number of components chosen automatically between 1 and 10 based on the Akaike Information Criterion (AIC). However, the number of components in the Gaussian mixture copula must be manually specified depending on the problem.
We compare our approach to a method using a Gaussian copula, referred to as GC. This method operates similarly to ours in learning the marginal distributions but assumes a Gaussian copula instead of a Gaussian mixture copula. The conditioning is performed in a latent Gaussian space.
Given the complexity of fitting a Gaussian Mixture Copula Model (GMCM) in the uniform space, an alternative approach for conditional sampling involves fitting a Gaussian Mixture Model (GMM) in a transformed space using the EM algorithm, which boasts theoretical convergence guarantees. Specifically, we can transform each marginal distribution to an approximate standard normal distribution. Following this, we can perform conditioning within this space and subsequently transform the samples back to the original space. We will refer to this method as Transformed Gaussian Mixture Model (TGMM).
Finally, we benchmark our approach against a Conditional Kernel Density Estimation (CKDE) method, where the joint density is estimated using KDE, and samples are drawn from the conditional density by filtering based on localization conditions, see details in the appendix.
5.1 Synthetic examples
5.1.1 Simulation scenarios
We consider two simulation scenarios where the response is univariate:
  • Gaussian mixture model:
    The data points are generated from a Gaussian mixture model in dimension two, with known parameters.
    Here, we take $(\mathbf{X}_{1},\mathbf{X}_{2})\sim\text{GMM}(\{\alpha_{i},\mu_{i},\sigma_{i}\}_{i=1}^{2})$ with $\alpha_{1}=0.3,~{}\alpha_{2}=0.7$ and
    $\mu_{1}=\begin{pmatrix}4\\ 2\end{pmatrix},~{}\mu_{2}=\begin{pmatrix}-2\\ 1\end{pmatrix}$ and $\Sigma_{1}=\begin{pmatrix}2&1\\ 1&1\end{pmatrix},~{}\Sigma_{2}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$.
    Samples from this multivariate distribution are represented on figure 5.
    In this case, we know analytically the form of the conditional distribution:
    $X_{1}|X_{2}=x_{2}\sim\tilde{\alpha}_{1}(x_{2})\mathcal{N}(4+(x_{2}-2),1)+\tilde{\alpha}_{2}(x_{2})\mathcal{N}(\frac{(x_{2}-1)}{2}-2,0.75)$
    with $\tilde{\alpha}_{1}(x_{2})=\frac{0.3\mathcal{N}(x_{2};2,1)}{0.3\mathcal{N}(x_{2};2,1)+0.7\mathcal{N}(x_{2};1,1)}$ and $\tilde{\alpha}_{2}(x_{2})=\frac{0.7\mathcal{N}(x_{2};1,1)}{0.3\mathcal{N}(x_{2};2,1)+0.7\mathcal{N}(x_{2};1,1)}$.
    In this test case, the data follows a Gaussian mixture dependence structure, and we know the specific number of components. As a result, we anticipate that our method GMCM, will perform very well.
  • Meta Gaussian mixture model:
    Let’s now consider another example with a univariate response, where the data has been simulated from a $2$-dimensional meta-Gaussian mixture model. This means that the copula corresponds to that of a Gaussian mixture model, while the marginal distributions can have various forms.
    We consider a Gaussian mixture copula model with the same parameters as in the previous example. The marginals are chosen to be standard normal distributed. An illustration of this multivariate distribution can be found in figure 7.
5.1.2 Results
Since we have considered synthetic examples where the response is univariate and the cumulative distribution function (cdf) of the conditional distribution has an analytical form, we can visually compare the cdf of the true conditional distribution with the cdf of the estimated conditional distribution produced by our method, at various new locations. Additionally, we compare the various methods using univariate scoring rules.
In figure 3 and 4, we represent the true cdf of the conditional distribution and the estimated cdf by our method using GMCM, at various conditioning locations. The true cdf is depicted with a blue line and the cdf estimated by our method by a red line.
Figure 3: Comparison of true conditional cdf and estimated conditional cdf at various locations. The data follows a Gaussian mixture model distribution.
Figure 4: Comparison of true conditional cdf and estimated conditional cdf at various locations. The data follows a meta-Gaussian mixture model distribution.
Based on these figures, it appears that our algorithm accurately recovers the conditional distributions when the joint distribution is generated from a Gaussian mixture model or from a meta Gaussian mixture model, when the number of components is known.
Additionally, we compare the various methods using univariate scoring rules.
Table 3: Comparison of various methods based on the mean of scoring rules over the test set for the various simulation scenarios.
GMM Meta GMM
CRPS LogS CRPS LogS
GC 3.141 3.055 0.875 2.056
GMCM 1.277 1.860 0.436 1.172
TGMM 1.279 1.889 0.440 1.177
CKDE 1.367 2.090 0.530 1.479
The results indicate that our method using Gaussian Mixture Copula Model outperforms the competitors for conditional sampling on simple examples.
Notice that another advantage of our method is that it provides estimates of the cdf $\hat{F}_{\mathbf{X_{1}}|\mathbf{X_{2}}=\mathbf{x_{2}}}(\mathbf{x_{1}})$ which are both smooth in $\mathbf{x_{2}}$ but also in $\mathbf{x_{1}}$. However, we need to fix the number of mixtures components beforehand, which would require some prior knowledge about the data. One potential solution is to select a large number of mixture components to enhance the modeling capacity of our model, thereby capturing more complex dependence structures. Nonetheless, this approach introduces greater challenges in parameter estimation. Consequently, the advantage gained from a more comprehensive modeling of the joint distribution is offset by the reduced accuracy in parameter estimation.
5.2 Real world examples
We now turn to real datasets with multivariate outputs and use multivariate scoring rules for evaluation.
5.2.1 Wine dataset
This dataset already presented in the introduction contains the amount of $13$ constituents found in each of the three types of wines and and consists of $178$ observations. We will set the number of components in our GMCM to 3. For our analysis, we will work with a $5$-dimensional problem where we try to predict the $3$ first variables from joint observations of the $5$ first variables. A graphical representation of this dataset can be found in figure 9.
The results, in table 4, show that the GMCM performs similarly than GC and TGMM. However, the CKDE benchmark seems to outperform all the other methods on this example. This could be attributed to the modeling of marginal distributions, for which we assume a Gaussian mixture form in our methodology, and seems not to be adequate in this example.
Table 4: Comparison of various methods based on the mean of scoring rules over the test set for the Wine dataset.
ES VS
GC 1.254 0.470
GMCM 1.221 0.489
TGMM 1.308 0.491
CKDE 1.207 0.466
5.2.2 Breast Cancer data Wisconsin
The data set provides data for $569$ patients on $30$ features of the cell nuclei obtained from a digitized image of a fine needle aspirate (FNA) of a breast mass. For each patient the cancer was diagnosed as malignant or benign. This data-set has been studied for classification or clustering tasks.
In this section, we will work again with a $5$ dimensional problem by predicting the $3$ first variables from joint observations of the $5$ first variables. A graphical representation of the data-set can be found in figure 11.
Given that the dataset appears to contain two clusters corresponding to malignant and benign diagnoses, we will set the number of components in our GMCM to 2. The results in table 5 demonstrate that our method outperforms the other methods for conditional sampling on this example.
Table 5: Comparison of various methods based on the mean of scoring rules over the test set for the breast cancer data Wisconsin.
ES VS
GC 1.896 0.848
GMCM 1.845 0.823
TGMM 1.854 0.826
CKDE 2.007 1.979
6 Conclusion and perspectives
In this work, we have proposed an interpretable method allowing easily estimating and sampling from conditional distributions based on estimated joint distributions belonging to a broad class way beyond the Gaussian case. We introduced the notion of trans-dimensional families of multivariate distributions that are stable by marginalization and conditioning. This class encompasses well-known families of multivariate distributions such as multivariate Gaussian distributions, Student t-distributions, and skew normal distributions. We demonstrated that this class extends to mixtures and transformations, providing a wide range of multivariate distributions that remain stable by conditioning.
Leveraging this stability property, we focused on distributions with tractable formulas for their conditional distributions. We developed an algorithm to estimate or sample from conditional distributions by modeling the joint distribution with a meta distribution. This approach uses copulas that are stable by conditioning with arbitrary marginals. While our method is highly versatile, for practical reasons, we concentrated in applications on the Gaussian copula mixture model. In turn, we empirically compared different approaches to the associated copula parameter fitting problem.
Numerical results indicate that our approach can accurately estimate conditional distributions in simple synthetic cases. On examples with real datasets, the performance of our method is more variable and sometimes lower than the benchmarks. However, in cases where the data exhibits a multi-modal nature, leveraging multiple components within the Gaussian mixture copulas framework can offer significant benefits. Furthermore our method is interpretable and preserves privacy, which is not the case of the other benchmarks. In particular, comparing different classes of imputation (as well as synthetic generation) methods both in terms of scoring approaches such as used here and of privacy indicators would be of great interest for future work.
Given the promising results of our method in the considered continuous framework, an important next step is to extend this methodology to handle mixed variables, which include both continuous and discrete components. Such an extension would be particularly relevant in medical applications where health outcomes of interest are often measured as categorical or dichotomous states. A possible route would be to to assume the discrete variables are discretizations of latent continuous variables as is common practice in latent class modelling frameworks.
Future research directions also include extending our method by implementing the algorithm for other dependence structures, such as mixtures of Student t-distributions and mixtures of skew normal distributions. Additionally, we aim to extend our work to accommodate missing values in the observed sample.
Finally, a deeper exploration of the theoretical properties of the proposed method could yield new insights and potentially pave the way for novel applications across various fields.
Code and datasets
The code used to estimate and sample from the conditional distributions is available in Python. The datasets are publicly available online.
Wine data: https://archive.ics.uci.edu/ml/machine-learning-databases/wine/wine.data
Breast Cancer data: https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic
Disclosure Statement
The authors report that there are no competing interests to declare.
Funding
This work has been funded by the Multidisciplinary Center for Infectious Diseases (MCID) from the University of Bern, grant number MA 14.
Appendix A Multivariate Student $t$-distribution
Definition A.1.
(Multivariate Student $t$-distribution)
If $\textbf{Y}\sim\mathcal{N}_{d}(0,\Sigma)$ and $u\sim\chi^{2}_{\nu}$ then $\frac{\textbf{Y}}{\sqrt{\frac{u}{\nu}}}=\textbf{X}-\mu$ with $\textbf{X}\sim t_{d}(\mu,\Sigma,\nu)$, where $\chi^{2}_{\nu}$ represents a chi-square distribution with $\nu$ degrees of freedom. We say that X follows a n-multivariate t-distribution with mean $\mu$, covariance matrix $\Sigma$ and $\nu$ degrees of freedom.
The multivariate Student $t$-distribution is stable by marginalization and conditioning as states the two following propositions [11]. In this appendix, we will use fixed, ordered sets of indices, with the first set containing the variables we wish to condition on the variables in the second set. These results can be readily extended to any arbitrary set of indices.
Proposition A.2.
(Marginal distribution) Let $\textbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})^{t}\sim t_{d}(\mu,\Sigma,\nu)$ with $\mu=(\mu_{1},\mu_{2})^{t}$, $\mu_{1}\in\mathbb{R}^{\ell}$ , $\mu_{2}\in\mathbb{R}^{m}$, $\ell+m=d$ and
$\Sigma=\begin{pmatrix}\Sigma_{1,1}&\Sigma_{1,2}\\ \Sigma_{2,1}&\Sigma_{2,2}\end{pmatrix}$
then the marginal distribution of $\mathbf{X_{2}}$ is given by:
$\mathbf{X}_{2}\sim t_{m}(\mu_{2},\Sigma_{22},\nu)$
The marginal distribution is again a multivariate Student t-distribution.
Proposition A.3.
(Conditional distribution)
With the same notation as in the proposition above, the conditional distribution of $\mathbf{X}_{1}$ given that $\mathbf{X}_{2}=\mathbf{x_{2}}$ is:
$\mathbf{X_{1}}|\mathbf{X}_{2}=\mathbf{x_{2}}\sim t_{\ell}\left(\mu_{1.2},\frac{\nu+d_{2}}{\nu+\ell}\Sigma_{11.2},\nu+\ell\right)$
with $\mu_{1.2}=\mu_{1}+\Sigma_{12}\Sigma_{22}^{-1}(\textbf{x}_{2}-\mu_{2})$ , $\Sigma_{11.2}=\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}$ and $d_{2}=(\textbf{X}_{2}-\mu)^{t}\Sigma_{22}^{-1}(\textbf{X}_{2}-\mu)$.
The conditional distribution is again a multivariate Student t-distribution but with another degrees of freedom. We say that the characteristic generator has been modified through conditioning.
Appendix B Multivariate unified skew distributions
Another class of multivariate distributions which are stable by conditioning and marginalization property are the multivariate unified skew normal and skew Student $t$-distributions. Unlike the Gaussian distribution, they allow for skewness.
Multivariate skew normal distributions can be defined in different ways by extending the univariate skew-normal density to the $d$-dimensional case [4]. Here we are mainly interested in the Multivariate Unified Skew Normal distribution, we will explain his stochastic construction and conditioning properties.
Let $U=\begin{bmatrix}\mathbf{U}_{0}\\ \mathbf{U}_{1}\end{bmatrix}\sim\mathcal{N}_{p+d}(0,\Omega^{*})$ with $\Omega^{*}=\begin{bmatrix}\Gamma&\Delta^{T}\\ \Delta&\bar{\Omega}\end{bmatrix}$, $\mathbf{Z}=(\mathbf{U_{1}}|\mathbf{U_{0}}+\gamma>0)$ , with $\gamma\in\mathbb{R}^{d}$, and let define $Y=\xi+\omega Z$ with $\xi\in\mathbb{R}^{d}$ and $\omega=\text{diag}(\omega_{1},\ldots,\omega_{d})>0$.
The density function of a variable of type $(X_{1}|X_{0}>c)$, evaluated at point $\mathbf{x}$, is most easily computed via the general relationship
$f(\mathbf{x})=\frac{f_{X_{1}}(\mathbf{x})\mathbb{P}(X_{0}>c|X_{1}=\mathbf{x})}{\mathbb{P}(X_{0}>c)}$
It can be shown that the density of $\mathbf{Y}$ at $\mathbf{x}\in\mathbb{R}^{d}$ is:
$f_{Y}(\mathbf{x})=\phi_{d}(\mathbf{x}-\xi;\Omega)\frac{\Phi_{p}(\gamma+\Delta^{T}\bar{\Omega}^{-1}\omega^{-1}(\mathbf{y}-\xi));\Gamma-\Delta^{T}\bar{\Omega}^{-1}\Delta)}{\Phi_{p}(\gamma;\Delta)}$
then we write $Y\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\bar{\omega}=\omega 1_{d}$ the vector of diagonal elements of $\omega$, and we say that $Y$ as a multivariate skew unified normal distribution [2].
The two following properties shows that the multivariate skew unified normal distribution is closed under marginalization and conditioning.
Proposition B.1.
(Marginal distribution)
Let $\mathbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})^{t}\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\mathbf{X_{1}}\in\mathbb{R}^{\ell}$ and $\mathbf{X_{2}}\in\mathbb{R}^{m}$ with $\ell+m=d$, $\xi=(\xi_{1},\xi_{2})^{t}$ with $\xi_{1}\in\mathbb{R}^{\ell},~{}\xi_{2}\in\mathbb{R}^{m}$ and $\bar{\omega}_{2}=\text{diag}(\omega_{\ell+1},\ldots,\omega_{d})\in\mathbb{R}^{m\times m}$ The marginal distribution of $\mathbf{X_{2}}$ is
$\mathbf{X}_{2}\sim SUN_{m,p}(\xi_{2},\gamma,\bar{\omega}_{2},\Omega^{*}_{2})$
where $\Omega^{*}_{2}=\begin{bmatrix}\Gamma&\Delta_{2}^{t}\\ \Delta_{2}&\bar{\Omega}_{22}\end{bmatrix}$.
Proposition B.2.
(Conditional distribution)
With the same notations as above, the conditional distribution of $\mathbf{X}_{1}$ given $\mathbf{X}_{2}=\mathbf{x}_{2}$ is given by
$(\mathbf{X}_{1}|\mathbf{X}_{2}=\mathbf{x}_{2})\sim SUN_{\ell,p}(\xi_{1.2},\gamma_{1.2},\bar{\omega}_{1},\Omega^{*}_{22.1})$
where $\xi_{1.2}=\xi_{1}+\Omega_{12}\Omega_{22}^{-1}(\mathbf{y}_{2}-\xi_{2})$, $\gamma_{1.2}=\gamma+\Delta_{2}^{T}\bar{\Omega}_{22}^{-1}\omega_{2}^{-1}-(\mathbf{y}_{2}-\xi_{2})$ and
$\Omega^{*}_{22.1}=\begin{bmatrix}\Gamma_{1.2}&\Delta_{1.2}^{T}\\ \Delta_{1.2}&\bar{\Omega}_{11.2}\end{bmatrix}$
with $\Omega^{*}_{11.2}=\bar{\Omega_{11}}-\bar{\Omega}_{12}\bar{\Omega}_{22}^{-1}\bar{\Omega}_{21}$, $\Delta_{1.2}=\Delta_{1}-\bar{\Omega}_{12}\bar{\Omega}_{22}^{-1}\Delta_{2}$ and $\Gamma_{1.2}=\Gamma-\Delta_{2}^{T}\bar{\Omega}_{22}^{-1}\Delta_{2}$.
With the stochastic representation of a multivariate skew unified normal distribution detailed above, we can generate samples from this distribution if we know how to sample from a multivariate normal distribution.
Appendix C Experiment details
C.1 GMCM fitting
For the AD implementation, we utilized the Adam optimizer with a learning rate of $10^{-3}$. For the FD method, the Nelder-Mead optimizer was employed, with a maximum of 10,000 iterations. Similarly, the PEM method also used a maximum of 10,000 iterations. The true values of the parameters are the following:
  • 2D example:
    $\alpha_{1}=0.45,~{}\alpha_{2}=0.55$, $\mu_{1}=\begin{pmatrix}5.15\\ 4.32\end{pmatrix},~{}\mu_{2}=\begin{pmatrix}-20.07\\ 3.04\end{pmatrix}$, $\Sigma_{1}=\begin{pmatrix}5.6&2.3\\ 2.3&8.27\end{pmatrix},~{}\Sigma_{2}=\begin{pmatrix}3.35&1\\ 1&1.16\end{pmatrix}$
  • 3D example:
    $\alpha_{1}=0.69,~{}\alpha_{2}=0.163,~{}\alpha_{2}=0.147$, $\mu_{1}=\begin{pmatrix}1.19\\ 5.63\\ -9.67\end{pmatrix},~{}\mu_{2}=\begin{pmatrix}-6.75\\ 12.04\\ -8.44\end{pmatrix},~{}\mu_{3}=\begin{pmatrix}-5.92\\ -4.0\\ 2.58\end{pmatrix}$, $\Sigma_{1}=\begin{pmatrix}2.26&1.33&-1.163544\\ 1.33&2.71&-1.36\\ -1.16&-1.36&3.78\par\end{pmatrix},~{}\Sigma_{2}=\begin{pmatrix}12.24&4.14&-3.87\\ 4.14&10.33&2.76\\ -3.87&2.76&16.45\end{pmatrix}$,
    $\Sigma_{3}=\begin{pmatrix}3.19&-0.38&0.54\\ -0.38&0.76&-0.02\\ 0.54&-0.02&1.01\end{pmatrix}$
C.2 CKDE Benchmark
The CKDE method used as benchmark, fits a Kernel Density Estimation (KDE) model (using a Gaussian kernel and a bandwidth of 1) on the joint distribution of the response variable $\mathbf{Y}$ and predictor $\mathbf{X}$. It then generates samples from the conditional distribution $\mathbb{P}(\mathbf{Y}|\mathbf{X})$. This is achieved by drawing an oversampled set of joint samples from the KDE model, filtering those to include only the samples close to a given predictor value $\mathbf{X}_{\text{cond}}$ within a specified tolerance (chosen to be 0.1) , and finally selecting the required number of samples (1000) from this filtered set. This process allows obtaining samples from the conditional distribution without needing an analytical form of the joint or conditional distributions.
C.3 Scoring rules
The CRPS, Energy Score and Variogram Score have been implemented using directly the Scoringrules package in Python. To implement the logarithmic score, we estimated the density of the predicted values by kernel density estimation (with a Gaussian kernel and a bandwidth of 0.5) and then evaluate the log-likelihood of the observed value.
Appendix D Additional figures
Figure 5: Visualization of the joint and marginal distributions of samples from a Gaussian Mixture Model.
Figure 6: Fitting of a GMCM on samples from a Gaussian Mixture Model.
Figure 7: Visualization of the joint and marginal distributions of samples from a meta Gaussian Mixture Model.
Figure 8: Fitting of a GMCM on samples from a meta-Gaussian Mixture Model.
Figure 9: Visualization of the joint, marginal and copulas distributions of the Wine data set.
Figure 10: Fitting of a GMCM on some components of the Wine data set.
Figure 11: Visualization of the joint, marginal and copulas distributions of the breast cancer data Wisconsin.Red points represent patients diagnosed with malignant, while green points represent patients diagnosed with benign.
Figure 12: Fitting of a GMCM on some components of the breast cancer data Wisconsin.