### v4

# Easy Conditioning far beyond Gaussian

Abstract
We address the challenge of conditioning multivariate densities, extending analytical conditioning results far beyond the Gaussian case. We review and discuss families of multivariate distributions that do enjoy analytical conditioning, also providing a few counter-examples. Proving that trans-dimensional stability under conditioning extends to mixtures and transformations, we demonstrate that a broader class of multivariate distributions inherit easy conditioning properties. Building on this insight, we developed a generative method to estimate conditional distributions from data by first fitting a flexible joint distribution using copulas and then performing analytical conditioning in a latent space. We specifically apply this methodology to Gaussian Mixture Copula Models (GMCM) and examine various fitting strategies. Through simulations and real-world data experiments, we showcase the efficacy of our method in tasks involving conditional density estimation and data imputation.
1 Introduction
Conditioning a multivariate density is a crucial yet challenging task in statistics. While conditional densities are easily obtained in the Gaussian case, more flexible multivariate distributions often require computationally intensive methods to obtain approximate conditional samples. In contrast, fast (approximate) evaluations of conditional probability density functions and cumulative distributions functions may be advantageous for computing quantities of interest, such as conditional quantiles and dependence measures.
In this paper, we explore non-Gaussian families of multivariate distributions that allow analytical computation of conditional distributions such as multivariate Student $t$-distribution [ding2016conditional] or multivariate skew normal distributions [azzalini1996multivariate]. Furthermore, by introducing the concept of trans-dimensional families of probability distributions that remain stable under conditioning, and showing that this principle extends to mixtures and transformations, we demonstrate that a broader class of distributions can benefit from analytical conditioning. In short, analytical conditioning of multivariate densities can be extended far beyond the Gaussian case.
This insight leads to a flexible generative method for estimating conditional distributions from data. The objective is to estimate the conditional distribution of a random vector $\mathbf{X_{1}}\in\mathbb{R}^{q}$ given an observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}\in\mathbb{R}^{p}$, based on a sample of moderate size and taking values in a continuous range, $(\mathbf{x_{1i}},\mathbf{x_{2i}})_{i=1}^{n}$ previously acquired. This setting is common in statistics, such as in probabilistic forecasting or imputation of systematically missing values, for example in medical or epidemiological applications.
Assuming conditions of existence to be fulfilled, a natural approach to estimate the conditional density $f(\mathbf{x_{1}}|\mathbf{x_{2}})$ is to first estimate the joint density by $\hat{f}(\mathbf{x_{1}},\mathbf{x_{2}})$ and then perform conditioning. Using our theoretical results, we show that this strategy can be applied far beyond the Gaussian case, which is often limiting due to its inability to capture complex dependencies and asymmetries typical in observed data. Notably, our approach allows modeling separately the dependence structure using copulas and the marginal distributions. Our method involves two main steps. First, we fit a copula to the data using a two-step procedure.
Next, we transform the original samples into a latent space, where conditioning is performed using analytical formulas and then we convert the samples back to the original space.
For applications, we will mainly focus on the Gaussian mixture copula model (GMCM) [tewari2011parametric] due to its ease of implementation and its ability to capture multi-modal dependencies. Additionally, we demonstrate that incorporating automatic differentiation enhances the parameter fitting process of GMCM compared to conventional approaches for computing derivatives [tewari2023estimation]. We conduct numerical experiments using both simulated and real data to demonstrate the relevance of our approach, evaluating the results with (multivariate) scoring rules.
As a motivating example, we first consider here the Wine data-set, which represents chemical analysis of wines grown in the same region in Italy. The analysis determined the quantities of 13 constituents found in each of the three types of wines. For illustrative purposes, we focus on $2$ constituents, the Alcohol level and Malic acids concentration. Figure 1 illustrates the approach proposed using these variables. The joint density appears here to be rather well modeled by a GMCM, the conditional distribution of which can be estimated and simulated from efficiently as detailed throughout the paper.
Figure 1: Steps of our conditional density estimation approach: from a sample of the joint distribution (left), we estimate the joint density by a mixture copular model (middle) and for for a given new value of Malic acids (5.04, red dotted line), we estimate (or sample from) the conditional distribution of Alcohol level by using analytical formulas (right).
The remainder of the paper is organized as follows. In Section 2, we introduce the properties of stability by marginalization and conditioning for families of multivariate distributions, provide some example and counter-examples families and show that these properties can be extended to mixtures and transformations. In Section 3, we present our method for estimating efficiently a conditional distribution by assuming a parametric form for the joint distribution and leveraging analytical formulas for conditioning. Section 4 focuses on the implementation of our algorithm for a particular case which is the GMCM. For the demonstration of the usefulness of our method, we perform numerical simulations and include real data examples in Section 5. Finally, in Section 6, we draw conclusions and outline areas for future research.
2 Stability by conditioning
2.1 Multivariate distributions stable by conditioning
2.1.1 Definition
Let us first consider a fixed dimension $d\geq 1$ and a product of spaces $\prod_{k=1}^{d}\mathcal{X}_{k}$. The $\mathcal{X}_{k}$ are assumed to be measure spaces, typically $\mathbb{R}$ or intervals, each endowed with the Lebesgue measure or restrictions thereof. We consider absolutely continuous probability distributions on $\prod_{k=1}^{d}\mathcal{X}_{k}$ with positive probability density functions $f^{(d)}(\cdot;\theta)$, where $\theta$ denotes a parameter. In this framework, we work indifferently with probability density functions (pdfs) or cumulative distribution functions (cdfs).
We denote as $\mathcal{F}^{(d)}=\{f^{(d)}(\cdot;\theta);\theta\in\Theta^{(d)}\}$ a parametric family of pdfs on $\prod_{k=1}^{d}\mathcal{X}_{k}$ with parameters belonging to a set $\Theta^{(d)}$.
We define a trans-dimensional family of pdfs by $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $1\leq D\leq+\infty$. The reason why we are concerned with this notion is because we will question the stability of such families by marginalization and conditioning with respect to some components.
Let us now introduce sets of indices that will be used for marginal distributions. For a given dimension $d>1$ the sets of $\ell$ strictly ordered indices are:
$I_{\ell}^{(d)}=\{\mathbf{i}=(i_{1},\ldots,i_{\ell}),~1\leq i_{1}<\ldots<i_{\ell}\leq d\},~\text{with}~1\leq\ \ell\leq d.$
For conditional distributions, we define a set of pairs of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ with elements $\mathbf{i}$ and $\mathbf{j}$ satisfying:
  • $\mathbf{i}=(i_{1},\ldots,i_{\ell})$ set of $\ell$ indices with $1\leq i_{1}<\ldots<i_{\ell}\leq d$, $1\leq\ell\leq d-1$,
  • $\mathbf{j}=(j_{1},\ldots,j_{m})$ set of $m$ indices with $1\leq j_{1}<\ldots<j_{m}\leq d$, $1\leq m\leq d$-$\ell$,
  • $\text{set}(\mathbf{i})\cap\text{set}(\mathbf{j})=\emptyset$ where the set of a vector is the set of all its components.
For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$, $1\leq\ell\leq d-1$, and $(\mathbf{i},\mathbf{j})\in I_{\ell,d-\ell}^{d}$, we define the function:
$f_{\mathbf{i}}:(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{i_{k}}\rightarrow\int_{\prod_{k=1}^{d-\ell}\mathcal{X}_{j_{k}}}f(x)dx_{j_{1}}\ldots dx_{j_{d-\ell}}$
which applies the marginalization of a $d$-dimensional density over $d$-$\ell$ variables. Let us note that since $\mathbf{j}$ has $d$-$\ell$ components, it is completely defined by $\mathbf{i}$.
To extend this operation to a trans-dimensional setting, we extend the marginalization operator as follows.
For $d\geq 1$ and $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we define :
$\text{Marg}(f^{(d)}(\cdot;\theta))=\bigcup\limits_{\ell=1}^{d-1}\{f_{\mathbf{i}}(\cdot;\theta),~\mathbf{i}\in I_{\ell}^{(d)}\}$
Now if we apply Marg to the entire family of distributions $\mathcal{F}^{(d)}$, we get:
$\text{Marg}(\mathcal{F}^{(d)})=\left\{\text{Marg}(f^{(d)}(\cdot;\theta)),\theta\in\Theta^{(d)}\right\}$
and finally for a trans-dimensional family of distributions $\mathcal{F}$ we set:
$\text{Marg}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\text{Marg}(\mathcal{F}^{(d)})$
We say that the trans-dimensional family of probability distributions $\mathcal{F}$ is stable by marginalization if it satisfies $\text{Marg}(\mathcal{F})\subset\mathcal{F}$.
Let us now consider conditioning. Since the pdfs considered here are positive, all considered conditional distributions are well-defined. For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$, a pair of sets of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and observed components $(x_{j_{1}},\ldots,x_{j_{m}})$, we denote:
$f_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}})):(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{i_{k}}\rightarrow\frac{f_{\left[\mathbf{i},\mathbf{j}\right]}(x_{i_{1}},\ldots,x_{i_{\ell}},x_{j_{1}},\ldots,x_{j_{m}})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}$
where $\left[\mathbf{i},\mathbf{j}\right]=(i_{1},\ldots,i_{\ell},j_{1},\ldots,j_{m})$.
In the same manner as above for marginalization, we define a conditioning operator taking a pdf in a parametric family as input and returning all possible conditional densities i.e for all pairs of sets of strictly ordered indices and every possible set of conditioning variables. For a function $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we have :
$\text{Cond}(f^{(d)}(\cdot,\theta))=\left\{f^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}});\theta),~(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d},~(x_{j_{1}},\ldots,x_{j_{m}})\in\prod_{k=1}^{m}\mathcal{X}_{i_{k}}\right\}$
By applying it to the trans-dimensional family of pdf $\mathcal{F}$ we get:
$\text{Cond}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\left\{\text{Cond}(f^{(d)}(\cdot;\theta)),~\theta\in\Theta^{(d)}\right\}$
We say that trans-dimensional family of probability distributions $\mathcal{F}$ is stable by conditioning if it satisfies $\text{Cond}(\mathcal{F})\subset\mathcal{F}$.
2.1.2 Examples
Stability by marginalization and conditioning can be demonstrated by showing that the marginal or conditional distributions remain within the same family of distributions, albeit in a lower dimension. Here, we will present several examples of distribution families that exhibit this property, as well as two counter-examples.
  • Gaussian distribution: For a fixed dimension $d\geq 1$, we consider
    $\mathcal{F}^{(d)}=\left\{f^{(d)}(x_{1},\ldots,x_{d};\mu,\Sigma)=\frac{\exp(-\frac{1}{2}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu))}{\sqrt{(2\pi)^{d}|\Sigma|}},\theta=(\mu,\Sigma)\in\Theta^{(d)}\right\}$
    where $\Theta^{(d)}=\mathbb{R}^{d}\times\text{PD}_{d}$ where $\text{PD}_{d}$ is the set of all symmetric positive definite matrices on $\mathbb{R}^{d\times d}$. The trans-dimensional family of Gaussian distributions is defined by: $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D=+\infty$. By properties of the multivariate Gaussian distribution, we know that associated marginal and conditional distributions remain Gaussian with parameters that can be computed analytically via well-known formulas. We have that $\text{Marg}(\mathcal{F})\subset\mathcal{F}$ and $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, and thus the trans-dimensional family of Gaussian distributions is stable by marginalization and by conditioning.
  • Student $t$-distribution: Another example is the family of multivariate Student t-distributions. Unlike the multivariate Gaussian distribution, it can model tail dependencies and it includes an additional “degrees of freedom” parameter, which controls the shape of the distribution, with higher values leading to distributions that resemble the multivariate Gaussian, while lower values result in heavier tails.
    It is defined as a scaled mixture of Gaussian distributions. In particular, if $\textbf{Y}\sim\mathcal{N}_{d}(0,\Sigma)$ and $u\sim\chi^{2}_{\nu}$ are independent then $\textbf{X}=\frac{\textbf{Y}}{\sqrt{\frac{u}{\nu}}}+\mu$ is such that $\textbf{X}\sim t_{d}(\mu,\Sigma,\nu)$, where $\chi^{2}_{\nu}$ represents a chi-square distribution with $\nu$ degrees of freedom. We say that X follows a multivariate $t$-distribution with mean $\mu$, covariance matrix $\Sigma$ and $\nu$ degrees of freedom.
    For $d\geq 1$, we denote the family of multivariate Student $t$-distribution as:
    $\mathcal{F}^{(d)}=\left\{f^{(d)}(\mathbf{x};\mu,\Sigma,\nu),\theta=(\mu,\Sigma,\nu)\in\Theta^{(d)}=\mathbb{R}^{d}\times\text{PD}_{d}\times\left[0,+\infty\right[\right\}$
    with $f^{(d)}(\mathbf{x};\mu,\Sigma,\nu)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})(\nu\pi)^{\frac{d}{2}}|\Sigma|^{\frac{1}{2}}}(1+\nu^{-1}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu))^{-\frac{(\nu+d)}{2}}$.
    The multivariate Student $t$-distribution is stable by marginalization and conditioning [ding2016conditional]; more details are included in the appendix.
  • Other examples:
    Other examples of families of multivariate distributions stable by marginalization and conditioning are the multivariate unified skew normal and skew Student $t$-distributions [azzalini2013skew]. The family of elliptically contoured distributions also enjoys these properties [cambanis1981theory]. However, as shown in [fang2018symmetric], conditioning an elliptical distribution can alter its generator, and a closed-form expression for the conditional distribution is generally not available. This limitation restricts its applicability in the remainder of this paper. On a different note, Dirichlet distributions are stable by conditioning, up to scalings ensuring that resulting measures are supported on the unit simplex (in appropriate dimensions) [fang2018symmetric].
  • Counter example 1: Student $t$-distribution with fixed degrees of freedom parameter.
    Let us now consider a simple counter-example of a family of multivariate distributions which is stable by marginalization but not stable by conditioning. For a given $d\geq 1$, consider the family of multivariate Student $t$-distributions with a fixed degree of freedom $\nu\in\left[0,+\infty\right[$ i.e
    $\mathcal{F}^{(d)}_{\nu}=\left\{f^{(d)}(\mathbf{x};\mu,\Sigma),\theta=(\mu,\Sigma)\in\Theta^{(d)}_{\nu}=\mathbb{R}^{d}\times\text{PD}_{d}\right\}$
    where $f^{(d)}(\mathbf{x};\mu,\Sigma)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})(\nu\pi)^{\frac{d}{2}}|\Sigma|^{\frac{1}{2}}}\left(1+\nu^{-1}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu)\right)^{-\frac{(\nu+d)}{2}}$ and the trans-dimensional family is defined by $\mathcal{F}_{\nu}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}_{\nu}$ with $D=+\infty$. Then, by properties of the multivariate Student $t$-distribution [ding2016conditional], this family is stable by marginalization. However, it is not stable by conditioning, because the resulting Student conditional distributions possess degrees of freedom different from $\nu$.
  • Counter example 2: $q$-Exponential distributions.
    A multivariate $q$-exponential distribution [li2023bayesian], denoted as $q\text{-ED}_{d}(\boldsymbol{\mu},\mathbf{C})$ for $(\boldsymbol{\mu},\mathbf{C})\in\mathbb{R}^{d}\times\text{PD}_{d}$, has the following density for $\mathbf{x}\in\mathbb{R}^{d}$ ($d\geq 1$ being fixed, here):
    $p(\mathbf{x}\mid\boldsymbol{\mu},\mathbf{C},q)=\frac{q}{2}(2\pi)^{-\frac{d}{2}}|\mathbf{C}|^{-\frac{1}{2}}r^{\frac{q}{2}-1}\exp\left(-\frac{r^{\frac{q}{2}}}{2}\right),\quad r(\mathbf{x})=(\mathbf{x}-\boldsymbol{\mu})^{\top}\mathbf{C}^{-1}(\mathbf{x}-\boldsymbol{\mu})$
    The counter-example presented now shows that for a given index $q$, the corresponding family of distributions is not stable by conditioning. From [li2023bayesian], we know that the $q$-exponential distribution is consistent in the sense defined by Kano [kano1994consistency]. By considering, $d=2$, $\mu=(0,0)^{t}$, $C=I_{2}$ and $\mathbf{x}=(x_{1},x_{2})^{t}$, we then have for $x_{1}\in\mathbb{R}$ :
    $p(x_{1})\propto(x_{1}^{2})^{{\frac{q}{2}}-1}\exp{\left(-\frac{1}{2}x_{1}^{q}\right)}$
    By assuming $p(x_{1})\neq 0$, the conditional distribution can then be written:
    $p(x_{2}|x_{1})\propto\frac{p(x_{1},x_{2})}{p(x_{1})}\propto\frac{(x_{1}^{2}+x_{2}^{2})^{(\frac{q}{2}-1)}}{(x_{1}^{2})^{(\frac{q}{2}-1)}}\exp{\left(-\frac{(x_{1}^{2}+x_{2}^{2})^{\frac{q}{2}}}{2}+\frac{(x_{1}^{2})^{\frac{q}{2}}}{2}\right)}$
    By taking $q=4$ and focusing on the exponential term, we get that the exponential term is:
    $\exp{\left(-\frac{x_{2}^{4}-2x_{1}^{2}x_{2}^{2}}{2}\right)}$
    If the conditional distribution were $q$-exponential (with $q=4$) it would be possible to write the exponential term as $\exp(-\frac{r^{2}}{2})$ with $r$ a quadratic form in $x_{2}$, which is not the case.
    We conclude that the conditional distribution is not $q$-exponential with $q=4$.
2.2 Main properties
2.2.1 Extension to finite mixtures
Consider again a parametrized family $\mathcal{F}^{(d)}=\{f^{(d)}(\cdot;\theta);\Theta\in\Theta^{(d)}\}$ of pdfs. We define
${\text{Mix}(\mathcal{F}^{(d)})=\left\{\sum_{k=1}^{K}\alpha_{k}f^{(d)}(\cdot;\theta_{k}),~K\geq 1,~\theta_{1},\ldots,\theta_{K}\in\Theta^{(d)},~\alpha_{1},\ldots,\alpha_{K}\geq 0~\text{s.t}~\sum_{k=1}^{K}\alpha_{k}=1\right\}}$
For a trans-dimensional family of distributions $\mathcal{F}$, we define its mixing set of distributions as the union over the mixing sets of the families with fixed dimensions i.e
$\text{Mix}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\text{Mix}(\mathcal{F}^{(d)})$
We now present results that extend the concepts of stability by marginalization and conditioning to finite mixtures.
Theorem 2.1.
If a parametrized family of trans-dimensional probability distributions $\mathcal{F}$ is stable by conditioning (resp. marginalization) then $\text{Mix}(\mathcal{F})$ is also stable by conditioning (resp. marginalization).
Proof of Theorem 2.1.
Let consider $f$ in $\text{Mix}(\mathcal{F})$. There exists $d\in\{1,\ldots,D\}$, $K\geq 1$, $\theta_{1},\ldots,\theta_{K}\in\Theta^{(d)}$, $\alpha_{1},\ldots,\alpha_{K}\geq 0~\text{s.t}~\sum_{k=1}^{K}\alpha_{k}=1$ such that:
$f(\cdot)=\sum_{k=1}^{K}\alpha_{k}f^{(d)}(\cdot;\theta_{k})$
For every pair of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and conditioning variables $(x_{j_{1}},\ldots,x_{j_{m}})\in\prod_{k=1}^{m}\mathcal{X}_{j_{k}}$, we have that $\forall(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{j_{k}}$:
$\begin{split}f_{\mathbf{i}|\mathbf{j}}((x_{i_{1}},\ldots,x_{i_{\ell}})|(x_{j_{1}},\ldots,x_{j_{m}}))&=\frac{f_{[\mathbf{i},\mathbf{j}]}(x_{i_{1}},\ldots,x_{i_{\ell}},x_{j_{1}},\ldots,x_{j_{m}})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}\\ &=\frac{1}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}\sum_{k=1}^{K}\alpha_{k}f^{(d)}_{[\mathbf{i},\mathbf{j}]}(x_{i_{1}},\ldots,x_{i_{\ell}},x_{j_{1}},\ldots,x_{j_{m}};\theta_{k})\\ &=\sum_{k=1}^{K}\frac{\alpha_{k}f^{(d)}_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}};\theta_{k})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}f^{(d)}_{\mathbf{i}|\mathbf{j}}((x_{i_{1}},\ldots,x_{i_{\ell}})|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\\ &=\sum_{k=1}^{K}\tilde{\alpha}_{k}(x_{j_{1}},\ldots,x_{j_{m}})f^{(d)}_{\mathbf{i}|\mathbf{j}}((x_{i_{1}},\ldots,x_{i_{\ell}})|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\end{split}$
where $\tilde{\alpha_{k}}(x_{j_{1}},\ldots,x_{j_{m}})=\frac{\alpha_{k}f^{(d)}_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}};\theta_{k})}{f_{\mathbf{j}}(x_{j_{1}},\ldots,x_{j_{m}})}$ which depends only on indices in $\mathbf{j}$.
Since $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, then for every $k\in\{1,\ldots,K\},~f^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\in\mathcal{F}$ and the conditional distribution $f_{\mathbf{i}|\mathbf{j}}$ is a mixture of elements of $\mathcal{F}$ with weights $\tilde{\alpha}_{k}$ so we can conclude that $f_{\mathbf{i}|\mathbf{j}}\in\text{Mix}(\mathcal{F})$. The proof for marginalization is analogous. ∎
This result can similarly be applied to cdfs. Note that if we can compute analytically the conditioned mixing weights $\tilde{\alpha}_{k}$ and if we can sample from the conditional distributions of each component $f_{\mathbf{i}|\mathbf{j}}(\cdot;\theta_{k})$, we can produce samples from the conditionals of the overarching mixture of distributions. Although the authors are not aware of prior formulations of Theorem 2.1, it its worth remarking that the formula for the conditional distribution of a mixture of distributions has in fact been utilized in diverse contexts [gilardi2002conditional, spycher2009multivariate].
2.2.2 Extension to transformations
We now work with parametric families of cdfs on $\prod_{k=1}^{d}\mathcal{X}_{k}$, similarly denoted denoted $\mathcal{F}^{d}=\left\{F^{(d)}(\cdot;\theta);\theta\in\Theta^{(d)}\right\}=\left\{F^{(d)}_{\theta};\theta\in\Theta^{(d)}\right\}$ as done with pdfs in the previous section. Similarly, we denote resulting trans-dimensional families as $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D\leq+\infty$.
For fixed $d\geq 1$ and $\mathcal{X}^{\prime}_{i}(i=1,\ldots,d$) measure spaces, we further define
$\Xi^{d}=\left\{\xi=\begin{pmatrix}\xi_{1}\\ \vdots\\ \xi_{d}\end{pmatrix}~\text{s.t}~\xi_{i}\in\mathcal{C}^{1}_{\uparrow}(\mathcal{X}_{i},\mathcal{X}^{\prime}_{i}),~i=1,\ldots,d\right\}$
where $\mathcal{C}^{1}_{\uparrow}(\mathcal{X}_{i},\mathcal{X^{\prime}}_{i})$ is the class of increasing $\mathcal{C}^{1}$ diffeomorphisms from $\mathcal{X}_{i}$ to $\mathcal{X}^{\prime}_{i}$. We introduce a “pushforward” operator, which takes as input an element of a parametric family of cdfs and returns all possible cdfs obtained by applying marginal $\mathcal{C}^{1}_{\uparrow}$ diffeomorphisms to the components, by:
$\Xi^{d}\#F^{(d)}_{\theta}=\left\{F^{(d)}_{\theta}\circ\xi^{-1};\xi\in\Xi^{d}\right\}$
Note that elements of $\Xi^{(d)}\#F^{(d)}_{\theta}$ are probability distributions on $\prod_{i=1}^{d}\mathcal{X^{\prime}}_{i}$ and that the obtained probability distributions are still absolutely continuous.
By applying the operator to the whole family of cdfs $\mathcal{F}^{(d)}$ we get:
$\Xi^{d}\#\mathcal{F}^{(d)}=\left\{\Xi^{d}\#F^{(d)}_{\theta};\theta\in\Theta^{(d)}\right\}$
Defining finally the class of functions with increasing marginal $\mathcal{C}^{1}$ diffeomorphisms as
$\Xi=\bigcup\limits_{d=1}^{D}\Xi^{d},$
we can now apply the pushforward operator to a trans-dimensional family of cdfs $\mathcal{F}$ by:
$\Xi\#\mathcal{F}=\bigcup\limits_{d=1}^{D}\Xi^{d}\#\mathcal{F}^{(d)}$
We call the trans-dimensional family $\Xi\#\mathcal{F}$ the pushforward of $\mathcal{F}$. In the next section, we relate this concept to copulas and meta families of distributions. At this point, we extend the stability of marginalization and conditioning results to transformations.
Theorem 2.2.
If a parametrized family of trans-dimensional probability distributions $\mathcal{F}$ is stable by conditioning (resp. marginalization) then $\Xi\#\mathcal{F}$ is also stable by conditioning (resp. marginalization).
Proof of Theorem 2.2.
: Let $F\in\Xi\#\mathcal{F}$. There exists $d\in\{1,\ldots,D\},~\theta\in\Theta^{(d)},~\xi\in\Xi^{d}$ such that $\forall(x_{1},\ldots,x_{d})\in\prod_{k=1}^{d}\mathcal{X}_{k}$:
$F(x_{1},\ldots,x_{d})=(F^{(d)}\circ\xi^{-1})(x_{1},\ldots,x_{d};\theta)=F^{(d)}(\xi_{1}^{-1}(x_{1}),\ldots,\xi_{d}^{-1}(x_{d});\theta)$
For every pair of strictly of ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and conditioning variables $(x_{j_{1}},\ldots,x_{j_{m}})\in\prod_{k=1}^{m}\mathcal{X}_{j_{k}}$, we have that $\forall(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{j_{k}}$:
$\begin{split}F_{\mathbf{i}|\mathbf{j}}(x_{i_{1}},\ldots,x_{i_{\ell}}|x_{j_{1}},\ldots,x_{j_{m}})&=F^{(d)}_{\mathbf{i}|\mathbf{j}}(\xi_{i_{1}}^{-1}(x_{i_{1}}),\ldots,\xi_{i_{\ell}}^{-1}(x_{i_{\ell}})|\xi_{j_{1}}^{-1}(x_{j_{1}}),\ldots,\xi_{j_{m}}^{-1}(x_{j_{m}});\theta)\\ &=F^{(d)}_{\mathbf{i}|\mathbf{j}}(z_{i_{1}},\ldots,z_{i_{\ell}}|z_{j_{1}},\ldots,z_{j_{m}};\theta)\end{split}$
with $z_{i_{k}}=\xi^{-1}_{i_{k}}(x_{i_{k}}),k=1,\ldots,\ell$ and $z_{j_{k}}=\xi^{-1}_{j_{k}}(x_{j_{k}}),k=1,\ldots,m$. Since $\mathcal{F}$ is stable by conditioning, then $F^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot;\theta)$ belongs to to $\mathcal{F}$. The proof is similar for marginalization. ∎
Note that we have only considered univariate transformations. However, for a fixed set of conditioning indices, it is worth mentioning that by following the same reasoning as in the proof, we can extend the property to blockwise bijective transformations.
3 Easy conditioning with fitted stable by conditioning distributions
We now focus on the task of conditioning an absolutely continuous random vector $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2})$ taking values in $\prod_{k=1}^{d}\mathcal{X}_{k}=\mathbb{R}^{d}$ on the second group of variables $\mathbf{X}_{2}$. On our approach is to first fit a joint (easy to condition) distribution to previously acquired sample $(\mathbf{x_{1i}},\mathbf{x_{2i}})_{i=1}^{n}$, and then, given any new observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}$, to leverage our results to estimate and/or sample from the conditional distribution of $\mathbf{X}_{1}$ given that $\mathbf{X}_{2}=\mathbf{x}_{2}$. Our generative procedure to estimate a conditional distribution thus consists in two main steps:
  • Fit an “easy to condition” joint distribution to a sample from $\mathbf{X}$
  • perform conditioning of the estimated joint by leveraging analytical formulas.
By using univariate transformations, we can model distributions with arbitrary (absolutely continuous) marginals. Specifically, if we consider the univariate transformations to be the marginal cumulative distribution functions, we obtain a space with uniform marginals and can subsequently apply copula theory, as detailed in the next section. In order to perform easy conditioning, we favor copula models corresponding to distributions stable by conditioning. Gaussian copulas, in particular, offer analytical formula for conditional distributions. However, they rely on latent multivariate Gaussian distributions, which are known to have notable limitations in terms of adequately model skewed, heavy-tailed, or multi-modal data. Alternative multivariate distributions, such as Student’s t-distributions [ding2016conditional] or skew-normal distributions [azzalini2013skew], offer more expressiveness and are stable by conditioning (see appendix for details). Resulting copula models can be used in our framework as they enable easy conditioning too. Yet, these distributions also fall short in scenarios requiring the modeling of multi-modal data and are confined to have marginal distributions of the same family along each axis. An important aspect of our the practical approach presented next is to allow capturing multi-modality by leveraging the extension of stability by conditioning to mixtures. The workhorse is to appeal to copulas of mixtures of multivariate distributions, where each individual component multivariate distribution possesses known conditional distributions.
3.1 Fitting joint distributions in terms of marginals and copulas
From a $d$-dimensional sample $\mathcal{D}=(x_{1i},\ldots,x_{di})_{i=1}^{n}$ of independently drawn realizations, we want to estimate the distribution of the underlying random vector $\mathbf{X}$ taking values in $\mathbb{R}^{d}$. We use here a two-step procedure based on copulas.
Copulas provides a versatile framework for multivariate analysis. Following Sklar’s theorem [sklar1959fonctions], for any distribution function $F$ with continuous margins $F_{j},~j=1,\ldots,d$, there exists a unique copula $C$ such that
$F(x_{1},\ldots,x_{d})=C(F_{1}(x_{1}),\ldots,F_{d}(x_{d})),~\forall\textbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$
A multivariate distribution function can then be modeled separately by its marginals and its copula where the copula characterizes the dependence structure between the variables.
There exist several parametric families of copulas. In particular, for a given absolutely continuous multivariate distribution, we denote by “implicit copula” the underlying the latter.
By “inverting” Sklar’s theorem, if we assume that the marginal distributions $F_{j},~j=1,\ldots,d$ are strictly increasing and continuous, we obtain that the copula can be written:
$C(u_{1},\ldots,u_{d})=F(F_{1}^{-1}(u_{1}),\ldots,F^{-1}_{d}(u_{d})),~0\leq u_{1},\ldots,u_{d}\leq 1$
For example, the Gaussian copula, which is the copula implicit in a multivariate Gaussian distribution can be expressed:
$C_{P}(\mathbf{u})=\Phi_{P}(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d}))$
where $\Phi$ is the cdf of the univariate standard normal distribution and $\Phi_{P}$ is the cdf of the multivariate $\mathcal{N}_{d}(\mathbf{0},\mathbf{P})$ distribution with $\mathbf{P}$ a correlation matrix.
Differentiating the general formula with respect to $\mathbf{u}$, we get the density of the implicit copula by:
$c(u_{1},\ldots,u_{d})=\frac{\partial^{d}}{\partial u_{1}\ldots\partial u_{d}}C(u_{1},\ldots,u_{d})=\frac{f(z_{1},\ldots,z_{d})}{\prod_{i=1}^{d}f_{i}(z_{i})}$
where $z_{j}=F_{j}^{-1}(u_{j}),~j\in\{1,\ldots,d\}$, we say that $\mathbf{z}=(z_{1},\ldots,z_{d})$ is the quantile transform of $\mathbf{u}$, and $f$ and $f_{i}$ denote the densities associated with $F$ and $F_{i}$, respectively, for all $i\in{1,\ldots,d}$.
If the marginals are known, we can transform the observations $\mathbf{x}_{i}$ to pseudo-observations $\mathbf{u}_{i}$, with uniform marginals, by using the Probability Integral Transform (PIT), $u_{ji}=F_{j}(x_{ji}),~\forall j\in\{1,\ldots,d\}$. And then to fit a parametric copula $c(u_{1},\ldots,u_{d};\theta)$ to the pseudo-observations by estimating the copula parameters $\theta$ with maximum-likelihood.
The optimization problem is then:
$\hat{\theta}\in\underset{\theta\in\Theta}{\mathrm{argmin}}\prod_{i=1}^{n}c(u_{1i},\ldots,u_{di};\theta)$
However, in practice the marginal distributions are typically unknown, and one needs to estimate them from the data. Parametric or non-parametric methods can be used to estimate the marginal distributions. Parametric estimation of the marginals is used in the inference for margins (IFM) method, while non-parametric estimation using the empirical distribution function is part of the pseudo likelihood estimator used in [genest1995semiparametric], which has been proven to be allow for asymptotically consistency in the estimation of dependence parameters.
Once the marginals have been estimated, then we can plug-in the “estimated pseudo observations” , $\hat{u}_{ji}=\hat{F}_{j}(x_{ji})$ in the maximum-likelihood optimization problem.
In our procedure, we assume that the joint distribution of the data is a meta distribution. A $d$-variate meta distribution is constructed by linking a $d$-variate implicit copula with $d$ arbitrary marginal cdfs. For example, a meta Gaussian distribution utilizes a Gaussian copula paired with arbitrary marginals.
In this study, we work with parametric copulas implicit in a family of distributions stable by conditioning. Consequently, as derived from Theorem 2.2, when considering strictly increasing marginal cdfs as a case of $\mathcal{C}_{\uparrow}^{1}(\mathbb{R},\mathbb{R)}$, the meta distributions we consider also exhibit stability by conditioning. For practical applications, our focus will be on copulas implicit in families of distributions for which analytical computation of conditional distributions is possible.
3.2 Conditioning via latent space
Once we have estimated the joint distribution, we estimate the conditional distribution (get conditional samples) given a new observation of the covariates. Since we assume that the data follow a "meta stable" distribution in order to perform easy conditioning, we are able to find a latent space in which we have stability by conditioning. In practice, we will focus on distributions for which there is an analytical formula for the conditional distributions.
Copulas transport the problem from the original space $\mathcal{X}$ to uniform space $\mathcal{U}=[0,1]^{d}$. We go one step further by transforming the marginals to a latent space $\mathcal{Z}$ [smith2023implicit], in which the marginals are the those of an implicit multivariate distribution driving the copula. The two consecutive transformations are illustrated in the following scheme:
[…]
By our assumption of "meta stable" distribution, we can use analytical formulas for conditioning in the latent space. Finally we transform back to the original space to get an estimate of the conditional distribution with the right marginals.
The full pseudo-algorithm to estimate a conditional distribution function is presented next:
Algorithm 1 Estimation of conditional CDF via latent space
[…]
  • Estimate the marginal distributions $F_{j},~j\in\{1,\ldots,d\}$ by $\hat{F}_{j}$.
  • Transform each marginal to approximately uniform by PIT, $\hat{u}_{j}=\hat{F}_{j}(x_{j})$.
  • Learn the copula function, which represents dependence structure by assuming a parametric form $c(\hat{u}_{1},\ldots,\hat{u}_{d},\theta)$. This is the implicit copula of a multivariate distribution with marginals $\Psi_{1},\ldots,\Psi_{d}$.
  • Transform the new observed point to the latent space $\mathcal{Z}$ by applying the transform
    $\mathbf{z}^{(new)}_{\ell+1:d}=(z^{(new)}_{\ell+1},\ldots,z^{(new)}_{d})=(\Psi^{-1}_{\ell+1}\circ\hat{F}_{\ell+1}(x_{\ell+1}^{(new)}),\ldots,\Psi^{-1}_{d}\circ\hat{F}_{d}(x_{d}^{(new)}))$
  • Perform conditioning in the latent space by using the analytical formulas to get an estimate of the cdf:
    $\hat{F}_{\mathbf{Z}_{1}|\mathbf{Z}_{2}=z^{(new)}_{\ell+1:d}}(z_{1}^{(new)},\ldots,z_{\ell}^{(new)}|z^{(new)}_{\ell+1:d})$
  • Transform back to get an estimate of the cdf in the original space by:
    $\hat{F}_{\mathbf{X}_{1}|\mathbf{X}_{2}}\left(x_{1},\ldots,x_{\ell}\big|\mathbf{x}_{\ell+1:d}^{(new)}\right)=\hat{F}_{\mathbf{Z}_{1}|\mathbf{Z}_{2}}\left(\hat{F}_{1}^{-1}\circ\Psi_{1}(z_{1}^{(new)}),\ldots,\hat{F}_{\ell}^{-1}\circ\Psi_{\ell}(z_{\ell}^{(new)})\big|\mathbf{z}_{\ell+1:d}^{(new)}\right)$
The algorithm above can be easily adapted to sample from the conditional distribution. The full workflow for estimating conditional distributions is illustrated on a $2$-dimensional example in figure 2.
Figure 2: Workflow of the conditioning algorithm on a $2$-dimensional example.
4 Implementation of the method
Our method’s versatility stems from its ability to choose any dependence structure among families of copulas stable by conditioning, which includes mixtures, and to model the joint distribution with arbitrary marginals. The primary challenge in our procedure lies in fitting a copula in the uniform space. This involves balancing the complexity of the dependence structure against the precision with which we can estimate the copula parameters. For this paper, we have opted to use a Gaussian mixture copula with a limited number of components. Usng further dependence structures such as Student’s $t$-mixture copulas or Skew-Normal mixture copulas represent exciting avenues for future research.
4.1 Gaussian Mixture Copula Model
The Gaussian mixture copula [tewari2011parametric] is the copula implicit in a Gaussian mixture model. It allows to capture multi-modality in the dependence structure. The density of a Gaussian mixture model (GMM) with $K$ components is given by:
$\psi(x_{1},\ldots,x_{d},\theta)=\sum_{k=1}^{K}\alpha_{j}\phi(x_{1},\ldots,x_{d};\mu_{k},\Sigma_{k})$
with $\alpha_{k}\geq 0,~\forall k\in\{1,\ldots,K\}$ such that $\sum_{k=1}^{K}\alpha_{k}=1$, the mixing weights and $\mu^{(k)},~\Sigma^{(k)}$ the mean and covariance matrix of the $k$-th component. $\theta=\{\alpha_{i},\mu_{i},\Sigma_{i}\}_{i=1}^{K}$ is the set of all parameters and $\phi$ the pdf of a Gaussian distribution.
Then from the equation 3.1, we have that the density of a Gaussian mixture copula with parameters $\theta$ is given by:
$c_{gmc}(u_{1},\ldots,u_{d};\theta)=\frac{\psi(\Psi^{-1}_{1}(u_{1}),\ldots,\Psi^{-1}_{d}(u_{d}))}{\prod_{i=1}^{d}\psi_{i}(\Psi_{i}^{-1}(u_{i}))},~0\leq u_{1},\ldots,u_{d}\leq 1$
where $\psi_{j}$ and $\Psi_{j}^{-1}$ denote the marginal density and inverse distribution function of the GMM along the $j$-th dimension .
Similarly, we can define the Student $t$-mixture copula [massing2021clustering] or the unified Skew-Normal mixture copula as others implicit copulas.
4.1.1 Fitting of a GMCM
Estimating parameters of a GMCM is a challenging task for several reasons as it has been discussed in [tewari2023estimation].
The first challenge is that copula invariance under marginal increasing transformations causes a lack of identifiability of the parameters.
This can be seen through the following proposition [tewari2023estimation]:
Proposition 1.
(Non identifiability of GMCM)
Let $\theta=\{\alpha_{i},\mu_{i},\Sigma_{i}\}_{i=1}^{K}$ and $\theta^{\prime}=\{\alpha_{i},A\mu_{i}+\mathbf{b},A^{t}\Sigma_{i}A\}_{i=1}^{K}$ two valid set of parameters with $\mathbf{b}\in\mathbb{R}^{d},~A\in\mathbb{R}^{d\times d}$. Then the likelihood of the two Gaussian mixture copula models with these parameters are the same i.e $\forall\mathbf{u}\in\mathbb{R}^{d},~c_{gmc}(u_{1},\ldots,u_{d};\theta)=c_{gmc}(u_{1},\ldots,u_{d};\theta^{\prime})$
This lack of identifiability results in the likelihood of a GMCM having infinitely many maximizers. To address this issue, Bilgrau [bilgrau2016gmcm] proposed standardizing the first component to have zero mean and unit variance along each dimension.
Another challenge in fitting GMCM is the high computational cost associated with functional evaluations, as the quantile function of a Gaussian mixture model isn’t known in closed form. To address this, we can use numerical approximation schemes leveraging the monotonicity of the cdf, although they tend to be computationally expensive. Alternatively, methods based on the bisection algorithm, such as the one developed by Chandrupatla [chandrupatla1997new], can be employed to find the inverse of a cdf. Since the derivatives of the Gaussian mixture model’s cdf are known analytically, a Newton-type algorithm can also be effectively used in this context.
In the literature, two primary methods have been proposed for estimating the parameters of a GMCM via direct gradient-based optimization. The classical approach involves using finite difference methods (denoted as FD in the rest of this paper) to estimate the derivatives. Approximating derivatives numerically can cause imprecisions and heavy computational costs.
Recently, a more efficient method, leveraging modern probabilistic programming language tools such as automatic differentiation (denoted as AD), has been developed. Automatic differentiation computes the exact values of the derivatives by leveraging the chain rule, significantly enhancing both computational efficiency and accuracy.
Another possible approach for parameter estimation is using an Expectation-Maximization (EM) type algorithm. However, due to the changing inverse distribution functions along the margins with every parameter update, the inputs of the EM algorithm are not fixed, rendering the traditional EM algorithm unsuitable. Although a pseudo-EM algorithm (denoted as PEM) has been proposed and discussed in the literature, it lacks convergence guarantees and often shows suboptimal performance.
The methods FD and PEM are implemented in the R package GMCM [bilgrau2016gmcm].Meanwhile, the AD method has been developed in Python by Tewari in a Github page related to this paper [tewari2023estimation].
4.1.2 Comparison of the copula fitting methods
To compare the different optimization methods described above (AD, FD and PEM), we simulate data following a GMCM with known parameters and attempted to recover its values using different fitting procedures. Because GMCM parameters are not identifiable, we cannot directly compare the recovered parameter values. Instead, we chose to compare metrics such as the log-likelihood or distances to the true distribution. We used the Cramér distance to quantify the differences between empirical multivariate distributions. We compared samples of the true underlying distributions to samples of the distribution with the fitted parameters and repeated the experiments several times to take into account the uncertainty in the sampling of a GMCM. In the results, we indicated the average Cramér distance and average log-likelihood as well as the standard deviations.
To ensure generality, we simulated data from several parameter configurations with varying dimensions and numbers of components, see details in the appendix.
Table 1: Fitting performances of the different optimization methods for a $2D$ example in terms of distance to a sample from the true underlying distribution.
Log-likelihood Cramer
Mean SD Mean SD
AD 208.7 16.9 0.278 $8.4\times 10^{-5}$
PEM -151.7 34.2 0.422 $9.1\times 10^{-5}$
FD -154.4 35.3 0.406 $8.8\times 10^{-5}$
Table 2: Fitting performances of the different optimization methods for a $3D$ example in terms of distance to a sample from the true underlying distribution.
Log-likelihood Cramer
Mean SD Mean SD
AD 1104.2 44.1 0.330 $8.9\times 10^{-5}$
PEM 1069.2 35.7 0.360 $8.80\times 10^{-5}$
FD 1063.8 43.7 0.364 $9.3\times 10^{-5}$
From the results above, it seems that the direct optimization of the likelihood function by using automatic differentiation tools (AD) performs better than the other methods for several parameter values. That is why we rely on this method for GMCM fitting in the experiments presented next.
5 Numerical experiments
We split the considered data sets randomly into a training set containing $80\%$ of the observations and a test set containing the remaining $20\%$ of the observations. We use the training set to learn the joint multivariate distribution and we sample $1000$ points from the estimated conditional distribution at every test points.
We then evaluate the performance of our approach comparing the conditional samples to the observed values using scoring rules. In the univariate case, we use the Continuous Ranked Probability Score (CRPS) and the logarithmic score. For multivariate outcomes, we employ the Energy score and the Variogram score [jordan2017evaluating], with order $r=0.5$. In particular, we will examine the mean value of the various scores across all test points, averaged over multiple train/test splits.
We denote the method presented in section 4 as GMCM. The marginal distributions are learned parametrically by assuming a Gaussian mixture model, with the number of components chosen automatically between 1 and 10 based on the Akaike Information Criterion (AIC). However, the number of components in the Gaussian mixture copula must be manually specified depending on the problem.
We compare our approach to a method using a Gaussian copula, referred to as GC. This method operates similarly to ours in learning the marginal distributions but assumes a Gaussian copula instead of a Gaussian mixture copula. The conditioning is performed in a latent Gaussian space.
Given the complexity of fitting a GMCM in the uniform space, an alternative approach for conditional sampling involves fitting a GMM in a transformed space using the EM algorithm, which boasts theoretical convergence guarantees. Specifically, we can transform each marginal distribution to an approximate standard normal distribution. Following this, we can perform conditioning within this space and subsequently transform the samples back to the original space. We will refer to this method as Transformed Gaussian Mixture Model (TGMM).
Finally, we benchmark our approach against a Conditional Kernel Density Estimation (CKDE) method, where the joint density is estimated using KDE, and samples are drawn from the conditional density by filtering based on localization conditions, see details in the appendix.
5.1 Synthetic examples
5.1.1 Simulation scenarios
We consider two simulation scenarios focused on sampling from a univariate conditional distribution:
  • Gaussian mixture model:
    The data points are generated from a Gaussian mixture model in dimension two, with known parameters.
    Here, we take $(\mathbf{X}_{1},\mathbf{X}_{2})\sim\text{GMM}(\{\alpha_{i},\mu_{i},\sigma_{i}\}_{i=1}^{2})$ with $\alpha_{1}=0.3,~\alpha_{2}=0.7$ and
    $\mu_{1}=\begin{pmatrix}4\\ 2\end{pmatrix},~\mu_{2}=\begin{pmatrix}-2\\ 1\end{pmatrix}$ and $\Sigma_{1}=\begin{pmatrix}2&1\\ 1&1\end{pmatrix},~\Sigma_{2}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$.
    Samples from this multivariate distribution are represented on figure 5.
    In this case, we know analytically the form of the conditional distribution:
    $X_{1}|X_{2}=x_{2}\sim\tilde{\alpha}_{1}(x_{2})\mathcal{N}(4+(x_{2}-2),1)+\tilde{\alpha}_{2}(x_{2})\mathcal{N}\left(\frac{(x_{2}-1)}{2}-2,0.75\right)$
    with $\tilde{\alpha}_{1}(x_{2})=\frac{0.3\mathcal{N}(x_{2};2,1)}{0.3\mathcal{N}(x_{2};2,1)+0.7\mathcal{N}(x_{2};1,1)}$ and $\tilde{\alpha}_{2}(x_{2})=\frac{0.7\mathcal{N}(x_{2};1,1)}{0.3\mathcal{N}(x_{2};2,1)+0.7\mathcal{N}(x_{2};1,1)}$.
    In this test case, the data follows a Gaussian mixture dependence structure, and we know the specific number of components.
  • Meta Gaussian mixture model:
    We now consider another example with a univariate response, where the data has been simulated from a $2$-dimensional meta-Gaussian mixture model. Here the marginal distributions are chosen standard normal, and the dependency structure is driven by a Gaussian mixture copula model with the parameters $\alpha_{1}=\alpha_{2}=0.5$ and
    $\mu_{1}=\begin{pmatrix}2\\ 2\end{pmatrix},~\mu_{2}=\begin{pmatrix}-2\\ -2\end{pmatrix}$ and $\Sigma_{1}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix},~\Sigma_{2}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$.
    An illustration of this multivariate distribution can be found in figure 7.
5.1.2 Results
Since we have considered synthetic examples where the response is univariate and the cumulative distribution function (cdf) of the conditional distribution has an analytical form, we can visually compare the cdf of the true conditional distribution with the cdf of the estimated conditional distribution produced by our method, at various new locations. Additionally, we compare the various methods using univariate scoring rules.
In figure 3 and 4, we represent the true cdf of the conditional distribution and the estimated cdf by our method using GMCM, at various conditioning locations. The true cdf is depicted with a blue line and the cdf estimated by our method by a red line.
Figure 3: Comparison of true conditional cdf and estimated conditional cdf at various locations. The data follows a Gaussian mixture model distribution.
Figure 4: Comparison of true conditional cdf and estimated conditional cdf at various locations. The data follows a meta-Gaussian mixture model distribution.
Based on these figures, it appears that our algorithm accurately recovers the conditional distributions when the joint distribution is generated from a Gaussian mixture model or from a meta Gaussian mixture model, when the number of components is known.
Additionally, we compare the various methods using univariate scoring rules.
Table 3: Comparison of various methods based on average scores over the test set for the various simulation scenarios.
GMM Meta GMM
CRPS LogS CRPS LogS
GC 1.207 1.878 0.328 0.922
GMCM 1.193 1.832 0.300 0.700
TGMM 1.214 1.864 0.318 0.870
CKDE 1.304 2.066 0.431 1.333
The results indicate that our method using GMCM and the TGMM outperform the other competitors for conditional sampling on simple examples.
In the next section, we turn to real datasets with multivariate outputs.
5.2 Real world examples
5.2.1 Wine dataset
This dataset already presented in the introduction contains the amount of $13$ constituents found in each of the three types of wines and consists of $178$ observations. We set the number of components in our GMCM to 3. For our analysis, we work with a $5$-dimensional problem where we try to predict the $3$ first variables given the $2$ remaining ones, based on training data consisting of from joint observations of the $5$ considered variables. A graphical representation of this dataset can be found in figure 9.
The results, in table 4, show that the GMCM performs better than GC and TGMM in terms of Energy Score (scoring rule in multivariate settings that generalizes CRPS). However, the CKDE benchmark seems to outperform all the other methods on this example.
Table 4: Comparison of various methods based on average scores over the test set for the Wine dataset.
ES VS
GC 1.263 0.504
GMCM 1.216 0.491
TGMM 1.290 0.493
CKDE 1.211 0.468
5.2.2 Breast Cancer data Wisconsin
The Breast Cancer Wisconsin dataset sourced from the UCI Machine Learning Repository, utilizes digitized images from fine needle aspirates (FNA) of breast masses. It comprises ten features derived from $569$ patients. For each feature, the dataset provides the mean value, extreme value (average of the three largest values), and standard error, resulting in $30$ continuous variables. The dataset includes two types of diagnoses: benign ($352$ patients) and malignant ($212$ patients), suggesting the presence of two distinct clusters. For this experiment, we focus on a subset of four variables: perimeter standard error (PSE), extreme smoothness (ES), extreme concavity (EC), and extreme concave points (ECP).
We aim to use probabilistic methods to predict the values of EC and ECP based on observations of PSE and ES.
A graphical representation of the data-set can be found in figure 11.
Given that the dataset appears to contain two clusters corresponding to malignant and benign diagnoses, we will set the number of components in our GMCM to 2. The results in table 5 demonstrate that our method and the TGMM perform competitively on this example.
Table 5: Comparison of various methods based on average scorings over the test set for the breast cancer data Wisconsin.
ES VS
GC 0.725 0.327
GMCM 0.671 0.285
TGMM 0.669 0.281
CKDE 1.012 0.523
5.3 Probabilistic imputation
One key advantage of the proposed approach is its flexibility: once the joint distribution is learned, we can analytically compute the conditional distribution for any combination of observed and missing variables This property is particularly valuable for imputation tasks, where different variables may be missing for different observations. In contrast, alternative distributional regression methods typically require fitting a separate model for each pattern of missing and observed variables.
Starting with a complete dataset, we randomly introduce missingness in a given proportion of entries (here, $10\%$). Observations with no missing values form the training set, while those with one or more missing values comprise the test set. We fit a Gaussian Mixture Copula Model (GMCM) to the training data. For each test observation, we sample from the conditional distribution of its missing variables, conditioned on its observed values. This produces multiple imputed datasets.
To evaluate imputation quality, we employ scoring rules. At first glance, applying scoring rules appears challenging: typically, one calculates an average score across all test predictions, but in our scenario, predictions may have different dimensions, complicating fair comparison.
One approach to circumvent this issue is to "complete" all predictions to joint distributions by plugging in the true observed values. However, it is worth questioning if evaluating these completed joint samples aligns with our actual imputation goal. Notably, for the Energy Score, it turns out that evaluating either the joint samples or the conditional samples of varying dimension is equivalent. In particular, for univariate predictions, this process reduces to calculating the CRPS.
We compared our approach with MICE [van2011mice], a widely used multiple-imputation procedure using chained equations. Our method was applied to a subset of four variables from the Breast Cancer Wisconsin dataset, see section 5.2.2, known for its multimodal structure, as well as a simulated dataset generated from a GMCM. For further details, we refer to Appendix C.3.
The results in Tables 7 and 6 indicate that GMCM slightly outperforms MICE in terms of Energy Score while requiring significantly less computational time. In experiments with a small number of features, our method performs comparably to MICE yet consistently shows a clear computational advantage. However, as the dimensionality increases, MICE tends to perform better. These findings suggest that our algorithm is well-suited for imputation tasks in datasets with few features and a multimodal structure, providing competitive results to MICE at a reduced computational cost.
Table 6: Imputation results for the Wisconsin data set
Metric GMCM MICE
Energy Score 0.252 0.286
Computational Time (s) 102.1 180.7
Table 7: Imputation results for simulated data
Metric GMCM MICE
Energy Score 0.097 0.102
Computational Time (s) 158.3 230.03
6 Conclusion and perspectives
In this work, we explored families of multivariate distributions that allow for analytical computation of conditional distributions. We provided examples beyond the Gaussian case, such as the multivariate Student $t$-distribution or the multivariate skew normal distribution.
Additionally, we introduced the concept of trans-dimensional families of multivariate distributions that are stable by conditioning. We demonstrated that this class extends to finite mixtures and transformations, showing that a broader range of multivariate distributions can benefit from analytical conditioning.
Leveraging these results, we developed a versatile approach for estimating conditional distributions from data. In practice, we model joint distributions using meta distributions with copulas that are stable by conditioning. In these settings, conditioning in original space is performed via a latent space where conditioning is more straightforward.
While our method allows for a high flexibility when considering broader families enjoying stability by conditioning, we focused in applications on the Gaussian Mixture Copula Model (GMCM) for practical purposes. In turn, we empirically compared different approaches to tackle the associated copula parameter fitting problem.
Numerical results indicate that our approach can accurately estimate conditional distributions in simple synthetic cases. Experiences on real datasets shows that the performance of our method is more variable and sometimes lower than the benchmarks. However, in cases where the data exhibits a multi-modal nature, leveraging multiple components within the GMCM framework can offer significant benefits. Moreover, our method offers several advantages over competitors: it is interpretable, provides an analytical form for the conditional distribution, and enables amortized inference for any combination of missing variables, which motivates its application to imputation tasks.
Given the promising results of our method in the considered continuous framework, an important next step is to extend this methodology to handle mixed variables, which include both continuous and discrete components. Such an extension would be particularly relevant in medical applications where health outcomes of interest are often measured as categorical or dichotomous states. A possible route would be to assume the discrete variables are discretizations of latent continuous variables as is common practice in latent class modeling frameworks. One limitation of our work is the challenge of fitting copulas, especially as dimensionality increases. Specifically, for the GMCM, the number of parameters scales quadratically with dimension, making it feasible only for relatively low-dimensional problems. Moreover, we currently need to manually preset the number of mixture components in the GMCM, whereas an automated method for determining this would be beneficial.
In higher-dimensional settings, other generative methods for sampling from a conditional density use tensor-train decomposition to approximate the joint density and sample from the conditional distribution by using the inverse Rosenblatt transform [cui2023scalable]. This approach has been shown to have strong theoretical guarantees, however it requires a predetermined ordering of variables, which can limit its usefulness in tasks like imputation.
Future research directions include extending our method by implementing the algorithm for other dependence structures, such as mixtures of Student t-distributions and mixtures of skew normal distributions. Additionally, we aim to extend our work to accommodate missing values in the observed sample.
Code and datasets
The code used in this paper has been developed in Python and is publicly available on GitHub. The datasets are publicly available online.
Funding
This work has been funded by the Multidisciplinary Center for Infectious Diseases (MCID) from the University of Bern, grant number MA 14.
Appendix A Multivariate Student $t$-distribution
Definition A.1.
(Multivariate Student $t$-distribution)
If $\textbf{Y}\sim\mathcal{N}_{d}(0,\Sigma)$ and $u\sim\chi^{2}_{\nu}$ then $\frac{\textbf{Y}}{\sqrt{\frac{u}{\nu}}}=\textbf{X}-\mu$ with $\textbf{X}\sim t_{d}(\mu,\Sigma,\nu)$, where $\chi^{2}_{\nu}$ represents a chi-square distribution with $\nu$ degrees of freedom. We say that X follows a n-multivariate t-distribution with mean $\mu$, covariance matrix $\Sigma$ and $\nu$ degrees of freedom.
The multivariate Student $t$-distribution is stable by marginalization and conditioning as states the two following propositions ([ding2016conditional]). In this appendix, we will use fixed, ordered sets of indices, with the first set containing the variables we wish to condition on the variables in the second set. These results can be readily extended to any arbitrary set of indices.
Proposition 2.
(Marginal distribution) Let $\textbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})^{t}\sim t_{d}(\mu,\Sigma,\nu)$ with $\mu=(\mu_{1},\mu_{2})^{t}$, $\mu_{1}\in\mathbb{R}^{\ell}$ , $\mu_{2}\in\mathbb{R}^{m}$, $\ell+m=d$ and
$\Sigma=\begin{pmatrix}\Sigma_{1,1}&\Sigma_{1,2}\\ \Sigma_{2,1}&\Sigma_{2,2}\end{pmatrix}$
then the marginal distribution of $\mathbf{X_{2}}$ is given by:
$\mathbf{X}_{2}\sim t_{m}(\mu_{2},\Sigma_{22},\nu)$
The marginal distribution is again a multivariate Student t-distribution.
Proposition 3.
(Conditional distribution)
With the same notation as in the proposition above, the conditional distribution of $\mathbf{X}_{1}$ given that $\mathbf{X}_{2}=\mathbf{x_{2}}$ is:
$\mathbf{X_{1}}|\mathbf{X}_{2}=\mathbf{x_{2}}\sim t_{\ell}\left(\mu_{1.2},\frac{\nu+d_{2}}{\nu+\ell}\Sigma_{11.2},\nu+\ell\right)$
with $\mu_{1.2}=\mu_{1}+\Sigma_{12}\Sigma_{22}^{-1}(\textbf{x}_{2}-\mu_{2})$ , $\Sigma_{11.2}=\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}$ and $d_{2}=(\textbf{X}_{2}-\mu)^{t}\Sigma_{22}^{-1}(\textbf{X}_{2}-\mu)$.
The conditional distribution is again a multivariate Student t-distribution but with another degrees of freedom. We say that the characteristic generator has been modified through conditioning.
Appendix B Multivariate unified skew distributions
Another class of multivariate distributions which are stable by conditioning and marginalization property are the multivariate unified skew normal and skew Student $t$-distributions. Unlike the Gaussian distribution, they allow for skewness.
Multivariate skew normal distributions can be defined in different ways by extending the univariate skew-normal density to the $d$-dimensional case ([azzalini1996multivariate]). Here we are mainly interested in the Multivariate Unified Skew Normal distribution, we will explain his stochastic construction and conditioning properties.
Let $U=\begin{bmatrix}\mathbf{U}_{0}\\ \mathbf{U}_{1}\end{bmatrix}\sim\mathcal{N}_{p+d}(0,\Omega^{*})$ with $\Omega^{*}=\begin{bmatrix}\Gamma&\Delta^{T}\\ \Delta&\bar{\Omega}\end{bmatrix}$, $\mathbf{Z}=(\mathbf{U_{1}}|\mathbf{U_{0}}+\gamma>0)$ , with $\gamma\in\mathbb{R}^{d}$, and let define $Y=\xi+\omega Z$ with $\xi\in\mathbb{R}^{d}$ and $\omega=\text{diag}(\omega_{1},\ldots,\omega_{d})>0$.
The density function of a variable of type $(X_{1}|X_{0}>c)$, evaluated at point $\mathbf{x}$, is most easily computed via the general relationship
$f(\mathbf{x})=\frac{f_{X_{1}}(\mathbf{x})\mathbb{P}(X_{0}>c|X_{1}=\mathbf{x})}{\mathbb{P}(X_{0}>c)}$
It can be shown that the density of $\mathbf{Y}$ at $\mathbf{x}\in\mathbb{R}^{d}$ is:
$f_{Y}(\mathbf{x})=\phi_{d}(\mathbf{x}-\xi;\Omega)\frac{\Phi_{p}(\gamma+\Delta^{T}\bar{\Omega}^{-1}\omega^{-1}(\mathbf{y}-\xi));\Gamma-\Delta^{T}\bar{\Omega}^{-1}\Delta)}{\Phi_{p}(\gamma;\Delta)}$
then we write $Y\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\bar{\omega}=\omega 1_{d}$ the vector of diagonal elements of $\omega$, and we say that $Y$ as a multivariate skew unified normal distribution ([arellano2006unification]).
The two following properties shows that the multivariate skew unified normal distribution is closed under marginalization and conditioning.
Proposition 4.
(Marginal distribution)
Let $\mathbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})^{t}\sim SUN_{d,p}(\xi,\gamma,\bar{\omega},\Omega^{*})$ with $\mathbf{X_{1}}\in\mathbb{R}^{\ell}$ and $\mathbf{X_{2}}\in\mathbb{R}^{m}$ with $\ell+m=d$, $\xi=(\xi_{1},\xi_{2})^{t}$ with $\xi_{1}\in\mathbb{R}^{\ell},~\xi_{2}\in\mathbb{R}^{m}$ and $\bar{\omega}_{2}=\text{diag}(\omega_{\ell+1},\ldots,\omega_{d})\in\mathbb{R}^{m\times m}$ . The marginal distribution of $\mathbf{X_{2}}$ is
$\mathbf{X}_{2}\sim SUN_{m,p}(\xi_{2},\gamma,\bar{\omega}_{2},\Omega^{*}_{2})$
where $\Omega^{*}_{2}=\begin{bmatrix}\Gamma&\Delta_{2}^{t}\\ \Delta_{2}&\bar{\Omega}_{22}\end{bmatrix}$.
Proposition 5.
(Conditional distribution)
With the same notations as above, the conditional distribution of $\mathbf{X}_{1}$ given $\mathbf{X}_{2}=\mathbf{x}_{2}$ is given by
$(\mathbf{X}_{1}|\mathbf{X}_{2}=\mathbf{x}_{2})\sim SUN_{\ell,p}(\xi_{1.2},\gamma_{1.2},\bar{\omega}_{1},\Omega^{*}_{22.1})$
where $\xi_{1.2}=\xi_{1}+\Omega_{12}\Omega_{22}^{-1}(\mathbf{y}_{2}-\xi_{2})$, $\gamma_{1.2}=\gamma+\Delta_{2}^{T}\bar{\Omega}_{22}^{-1}\omega_{2}^{-1}-(\mathbf{y}_{2}-\xi_{2})$ and
$\Omega^{*}_{22.1}=\begin{bmatrix}\Gamma_{1.2}&\Delta_{1.2}^{T}\\ \Delta_{1.2}&\bar{\Omega}_{11.2}\end{bmatrix}$
with $\Omega^{*}_{11.2}=\bar{\Omega_{11}}-\bar{\Omega}_{12}\bar{\Omega}_{22}^{-1}\bar{\Omega}_{21}$, $\Delta_{1.2}=\Delta_{1}-\bar{\Omega}_{12}\bar{\Omega}_{22}^{-1}\Delta_{2}$ and $\Gamma_{1.2}=\Gamma-\Delta_{2}^{T}\bar{\Omega}_{22}^{-1}\Delta_{2}$.
With the stochastic representation of a multivariate skew unified normal distribution detailed above, we can generate samples from this distribution if we know how to sample from a multivariate normal distribution.
Appendix C Experiment details
C.1 GMCM fitting
For the AD implementation, we utilized the Adam optimizer with a learning rate of $10^{-3}$. For the FD method, the Nelder-Mead optimizer was employed, with a maximum of 10,000 iterations. Similarly, the PEM method also used a maximum of 10,000 iterations. The true values of the parameters are the following:
  • 2D example:
    $\alpha_{1}=0.45,~\alpha_{2}=0.55$, $\mu_{1}=\begin{pmatrix}5.15\\ 4.32\end{pmatrix},~\mu_{2}=\begin{pmatrix}-20.07\\ 3.04\end{pmatrix}$, $\Sigma_{1}=\begin{pmatrix}5.6&2.3\\ 2.3&8.27\end{pmatrix},~\Sigma_{2}=\begin{pmatrix}3.35&1\\ 1&1.16\end{pmatrix}$
  • 3D example:
    $\alpha_{1}=0.69,~\alpha_{2}=0.163,~\alpha_{2}=0.147$, $\mu_{1}=\begin{pmatrix}1.19\\ 5.63\\ -9.67\end{pmatrix},~\mu_{2}=\begin{pmatrix}-6.75\\ 12.04\\ -8.44\end{pmatrix},~\mu_{3}=\begin{pmatrix}-5.92\\ -4.0\\ 2.58\end{pmatrix}$, $\Sigma_{1}=\begin{pmatrix}2.26&1.33&-1.163544\\ 1.33&2.71&-1.36\\ -1.16&-1.36&3.78\par\end{pmatrix},~\Sigma_{2}=\begin{pmatrix}12.24&4.14&-3.87\\ 4.14&10.33&2.76\\ -3.87&2.76&16.45\end{pmatrix}$,
    $\Sigma_{3}=\begin{pmatrix}3.19&-0.38&0.54\\ -0.38&0.76&-0.02\\ 0.54&-0.02&1.01\end{pmatrix}$
C.2 CKDE Benchmark
The CKDE method used as benchmark, fits a Kernel Density Estimation (KDE) model (using a Gaussian kernel and a bandwidth of 1 by using the function KernelDensity of the package sklearn in Python) on the joint distribution of the response variables and predictors $\mathbf{X}=(\mathbf{X_{1}},\mathbf{X_{2}})$ . It then generates samples from the conditional distribution $\mathbb{P}(\mathbf{X_{1}}|\mathbf{X_{2}})$. This is achieved by drawing an oversampled set of joint samples from the KDE model, filtering those to include only the samples close to a given predictor value $\mathbf{X}_{\text{cond}}$ within a specified tolerance (chosen to be 0.4 for most of the experiments, except for the experiment on the Wine dataset where it is chosen to be 10), and finally selecting the required number of samples (1000) from this filtered set. This process allows obtaining samples from the conditional distribution without needing an analytical form of the joint or conditional distributions.
C.3 Imputation experiment
The dataset for the imputation experiment on simulated data 5.3 has been generated by sampling $1000$ points from a Gaussian Mixture Copula Model (GMCM) with two components, with means $\mu_{1}=(1,2)^{t}$ and $\mu_{2}=(3,4)^{t}$, weights $w=(0.6,0.4)$ and covariances matrices:
$\Sigma_{1}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix},~~\Sigma_{2}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$
The marginal distributions have been chosen to be from a GMM with respectively means $\mu_{m,1}=(1,2)^{t}$ and $\mu_{2}=(3,4)^{t}$, weights $w_{1}=(0,4,0.6)$ and $w_{2}=(0.4,0.6)$, and standard deviations $\sigma_{1}=(1,1)^{t}$ and $\sigma_{2}=(1,1)^{t}$.
For the experiment with the Breast Cancer Wisconsin data, we focused on a subset of four variables: perimeter standard error (PSE), extreme smoothness (ES), extreme concavity (EC), and extreme concave points (ECP).
In both these experiments, we used the function qmixnorm from the R package KScorrect to compute quantiles of a Gaussian Mixture Model.
The computer used for this experiment is a HP Elite Book 845 with a a AMD Ryzen 7 6800U processor and 16 Go of RAM.
C.4 Scoring rules
The CRPS, Energy Score and Variogram Score have been implemented using directly the Scoringrules package in Python. To implement the logarithmic score, we estimated the density of the predicted values by kernel density estimation (with a Gaussian kernel and a bandwidth of 0.5) and then evaluate the log-likelihood of the observed value.
Appendix D Additional figures
Figure 5: Visualization of the joint and marginal distributions of samples from a Gaussian Mixture Model.
Figure 6: Fitting of a GMCM on samples from a Gaussian Mixture Model.
Figure 7: Visualization of the joint and marginal distributions of samples from a meta Gaussian Mixture Model.
Figure 8: Fitting of a GMCM on samples from a meta-Gaussian Mixture Model.
Figure 9: Visualization of all univariate and bivariate marginal combinations in both original and uniform spaces of the Wine data set.
Figure 10: Fitting of a GMCM on some components of the Wine data set.
Figure 11: Visualization of all univariate and bivariate marginal combinations in both original and uniform spaces for Wisconsin breast cancer data. Red points denote patients with malignant tumors, while green points represent those with benign tumors.
Figure 12: Fitting of a GMCM on some components of the breast cancer data Wisconsin.